### v2

# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

Abstract
In this work, we investigate the power-expansion/OPE expansion in super-renormalizable QFTs. In comparison to “marginal” short distance asymptotics, the power-expansion in super-renormalizable theory has a higher level of correlation with the coupling constant expansion, and the fixed-point perturbation theory for the coefficient functions suffers increasing level of IR divergences when the order of power-expansion/coupling constant expansion increases. This increasing IR sensitivity anticipates the presence of vacuum condensates of non-trivial local operators in the OPE expansion, which not only cancel the IR divergences for coefficient functions through their UV divergences, but also shape the structure of logarithms at each power in the power-expansion. This UV-IR correspondence has been used to argue for the renormalon cancellation between coefficient functions and operators in marginal asymptotics, where the residue IR sensitivity of coefficient functions to be canceled by the operators manifests through high-order factorial growths after renormalization subtraction.
However, in supper-renomalizable theory, high-order behavior in coupling constants in the coefficient functions is simultaneously large-order behavior in the power expansion, and one may wonder if there are also factorial growths for high-order/power terms in the coefficient functions, as they also require large-numbers of IR subtractions. In this work, we address the issue regarding the asymptotic behavior of power-expansion/OPE-expansion in supper-renormalizable theory. Using an $O(N)$-model with $N$-components scalars coupled through quartic interaction at the next-to-leading $\frac{1}{N}$ order in the large-$N$ expansion, we show that the IR subtractions indeed cause addition factorial enhancements for large order/power terms in the coefficient functions. Moreover, there are also factorial enhancements for the operator condensates, and the factorial enhancements cancel between coefficient functions and operators only off-diagonally across different powers. The factorial enhancements can be both alternating and non-alternating. The former is similar to the “UV renormalon” of coefficient functions and cancels with factorial enhancements of operators at lower powers in diagrams with negative degrees of UV divergences. The latter is similar to “IR renormalon” and cancels with factorial enhancements of renormalized operators at higher powers in diagrams with positive degrees of UV divergences. The factorial enhancement itself will render the momentum-space power expansion divergent.
1 Introduction
Operator product expansion (OPE) Wilson (1971) is among one of the major properties concerning short distance behavior of correlation functions in local QFTs. In the “simplest” local QFTs, the conformal field theories (CFTs), the existence of OPE in the sense of asymptotic expansion is sufficient to imply the absolute convergence of OPE in the coordinate-space analyticity region Luscher and Mack (1975); Mack (1977); Kravchuk et al. (2021). In massive QFTs with CFTs as their short-distance fixed-points, on the other hand, although still exist at worst in the sense of asymptotic expansion in the Euclidean region, the behavior of OPE is much less controlled due to the existence of non-trivial IR physics. Not only derivative operators are much harder to be resumed, the coefficient functions for individual operators are no longer homogeneous functions and will contain non-trivial $z^{2}m^{2}$ dependencies Novikov et al. (1985); Lukyanov and Zamolodchikov (1997) that couple the short distance and large distance scales.
In super-renormalizable QFTs, namely, QFTs that are obtained by perturbing the short-distance fixed points through relevant operators, the $z^{2}m^{2}$ dependencies in the coefficient functions can still be separated into well-defined powers and calculated order by order in fixed-point perturbation theory (also called CFT perturbation theory) with appropriate IR subtractions Chetyrkin et al. (1982); Chetyrkin (1983). The IR subtractions introduce scale-scheme dependencies to the coefficient functions that are canceled power-wisely in the OPE expansion, by the scale-scheme dependencies introduced to the operators during their UV-divergence renormalization. To some sense, the IR sensitivity in the fixed-point perturbation theory anticipates the existence of non-trivial operator condensates Novikov et al. (1980); Shifman et al. (1979); David (1984, 1986); Novikov et al. (1985); Mueller (1985); Ji (1995); Beneke (1999); Beneke et al. (1998); Braun et al. (2004); Shifman (2015); Dunne et al. (2015). However, exactly due to the large numbers of IR subtractions required to define the high-order terms in the coefficient functions, their quantitative behaviors can be significantly modified in comparison to quantities defined through absolutely convergent Feynman integrals. In the worst scenario, factorial growths can be developed at a relatively small number of diagrams and modify the convergence property of the OPE expansion or at least, the convergence of power-expansion for the coefficient functions.
It is the task of this work to investigate the issue of asymptotic behavior of coefficient functions in super-renormalizable QFTs in the presence of large numbers of IR subtractions. Using the large-$N$ expansion of an $O(N)$ model with quartic interaction, one can remove the factorial growths in the numbers of Feynman diagrams and focus on the renormalization effects. We show that the IR-subtraction-introduced factorial growths in the power expansion in the Euclidean momentum space indeed exist and are canceled between coefficient functions and operators off-diagonally across different powers. The off-diagonal nature of the cancellation implies that the power expansion in momentum space is divergent. To make the OPE expansion convergent, one must subtract all the factorial growths from operators and add back to coefficient functions. This requires infinitely many redefinitions of operators and can not be achieved through a single “Borel prescription”. On the other hand, in the coordinate space, due to the overall $\frac{1}{(n!)^{2}}$ suppression, the $n!$ introduced by the IR subtraction is insufficient to modify the convergence properties of the power-expansion/OPE expansion.
The paper is organized as follows.
  1. 1.
    In the Sec. 2, we provide a worked example of the OPE calculation up to all the non-trivial coefficient functions up to ${\cal O}(z^{4})$, for the scalar two-point function in the 2D one-component $\phi^{4}$ theory. The section is largely for pedagogical purposes to convince unfamiliar readers the fact that the coefficient functions in super-renormalized QFTs can still be calculated through IR-renormalized massless Feynman integrals regulated in dimensional regularization (DR). It will also demonstrate how the UV-IR correspondence between coefficient functions and operators works in this example.
  2. 2.
    In the Sec. 3, we start to investigate the asymptotic behavior of IR renormalized coefficient functions in the large-$N$ model. We first perform the calculation through IR renormalized massless bubble-chain diagrams, for the diagrams with zero and one mass insertions (but with $n$ and $n-1$ massless bubbles). The $n!$ enhancement is generated through a novel $\frac{1}{\epsilon}\times\epsilon$ effect that is absent in 4D bubble-chain calculations. We also show how the $n!$ enhancements cancel with that of operators in a way without directly calculating the massive bubble-chains. We then double-check with the short-distance calculation by introducing Mellin-Barnes representations for the full massive bubble chain diagram and performing the power-expansion by shifting the contours for the Mellin-Barnes. The result completely agrees with the short-distance calculation and allows one to effectively obtain the leading large-$n$ asymptotics for the coefficient function at the power $\frac{1}{(p^{2})^{n+2}}$ after summing over all possible numbers of mass insertions at the given power. Our work therefore provides a concrete example showing the “OPE divergence” phenomenon Shifman (1994) as asked in Marino et al. (2023).
  3. 3.
    Finally, we comment and conclude in Sec. 4.
2 OPE in supper-renormalizable theory: $\phi^{4}_{2}$ as an example
In this section, we provide a worked example for OPE calculation both in coordinate and momentum spaces, in the massive $\phi^{4}_{2}$ theory for the scalar two-point function. We always stay in Euclidean signature in this work. Although this section may look familiar to experts, the author still suggests reading this section, which could also serve as a motivation for later sections.
As a reminder, the action of the theory is given by
$S=\mu_{0}^{D-2}\int d^{D}x\bigg{(}\frac{1}{2}(\nabla\phi)^{2}+\frac{m^{2}}{2}\phi^{2}+\frac{g}{4!}:\phi^{4}:\bigg{)}\ .$ (1)
Throughout the paper we use DR with $D=d=2-2\epsilon$ for all the UV/IR divergences and $\mu_{0}^{2}=\mu^{2}\frac{e^{\gamma_{E}}}{4\pi}$ are the standard renormalization scales. Notice that the Wick-ordering is with respect to the free Gaussian field with mass $m$. This defines the mass parameter $m$, as well as the bare mass
$m_{0}^{2}=m^{2}-\frac{g\mu_{0}^{2-D}}{2}\int\frac{d^{D}k}{(2\pi)^{D}}\frac{1}{k^{2}+m^{2}}\equiv m^{2}-\frac{g}{2}I_{0}\ .$ (2)
In the $\epsilon\rightarrow 0$ limit, all correlation functions of the $\phi$ field at separated points are finite and can be calculated directly at $D=2$ by neglecting the one-loop tadpoles. Notice that the parameters $m$ and $g$ are $\epsilon$ independent.
The theory is well-defined and serves as the most general parameterization of 2D one-component stable quartic scalar field theory with $Z_{2}$ symmetry. In particular, one can set $m^{2}>0$ without losing of generality. This is because for $m^{2}<0$, the Wick-ordering must be performed with another mass $M^{2}>0$ without destroying reality, and one can always find $\tilde{m}^{2}>0$ such that
$m_{0}^{2}=m^{2}-\frac{g}{8\pi}\ln\frac{\mu^{2}}{M^{2}}-\frac{g}{8\pi\epsilon}+{\cal O}(\epsilon)=\tilde{m}^{2}-\frac{g}{8\pi}\ln\frac{\mu^{2}}{\tilde{m}^{2}}-\frac{g}{8\pi\epsilon}+{\cal O}(\epsilon)\ ,$ (3)
as far as $M^{2}>0$ even when $m^{2}<0$. The above implies that in the $\epsilon\rightarrow 0$ limit, the theory is equivalent to the standard one with $\tilde{m}^{2}>0$ at least order by order in the $\frac{g}{\tilde{m}^{2}}$ expansion. It is normally conjectured Chang (1976) that, when $g$ increase to $\frac{g}{m^{2}}=\frac{g_{c}}{m^{2}}>0$, there is a phase transition in the 2D Ising universality class from $Z_{2}$ unbroken to $Z_{2}$ broken phase, since $m_{0}^{2}$ becomes “too negative” at large $g$. The large distance behavior of the theory will be very different in the two phases, but if one is interested in the small distance behavior at $z^{2}g\ll 1,z^{2}m^{2}\ll 1$, then the two phases should share the same structure of the OPE expansion with identical small distance asymptotics power by power for the coefficient functions.
Below we show how the OPE calculation using DR regulated massless integrals works for the two point function
$S\left(z^{2}m^{2},\frac{g}{m^{2}}\right)=\langle\phi(z)\phi(0)\rangle=\int\frac{d^{2}p}{(2\pi)^{2}}e^{ip\cdot z}S\left(p^{2}\right)\ ,$ (4)
in this theory, up to all the coefficient functions contributing to ${\cal O}(z^{4})$. Since in the coefficient functions there are no inverse powers in $m^{2}$ and $g$, only $g^{2}$ (two-loop) is required for this calculation. To show that the OPE is correct, it is helpful to show that the coefficient functions combined with the operator condensates reproduce the correct small-$z^{2}$ expansion of tree-level and two-loop massive diagrams of $S$. In the appendix A, we show using explicit calculations that for the tree-level and two-loop two-point function
$S\left(z^{2}m^{2},\frac{g}{m^{2}}\right)=\int\frac{d^{2}p}{(2\pi)^{2}}\frac{e^{ip\cdot z}}{(p^{2}+m^{2})}+\frac{g^{2}}{6}\int\frac{d^{2}p}{(2\pi)^{2}}\frac{e^{ip\cdot z}}{(p^{2}+m^{2})^{2}}\Sigma(p^{2})\ ,$ (5)
$\Sigma(p^{2})=\int\frac{d^{2}kd^{2}l}{(2\pi)^{4}}\frac{1}{(k^{2}+m^{2})(l^{2}+m^{2})((p-k-l)^{2}+m^{2})}\ ,$ (6)
the $z^{2}m^{2}$ expansion up to $(z^{2}m^{2})^{2}$ reads
$S\left(z^{2}m^{2},\frac{g}{m^{2}}\right)=$
$-\frac{1}{4\pi}L+\frac{7\zeta_{3}}{24(4\pi)^{3}}\frac{g^{2}}{m^{4}}$
$+\bigg{(}\frac{2-L}{4\pi}-\frac{7\zeta_{3}}{8(4\pi)^{3}}\frac{g^{2}}{m^{4}}\bigg{)}\frac{z^{2}m^{2}}{4}$
$+\bigg{(}\frac{3-L}{16\pi}+\frac{g^{2}}{6(4\pi)^{3}m^{4}}\left(-\frac{L^{3}}{4}+\frac{9L^{2}}{4}-\frac{69L}{8}+\frac{27}{2}-\frac{49\zeta_{3}}{16}\right)\bigg{)}\left(\frac{z^{2}m^{2}}{4}\right)^{2}+{\cal O}(z^{6}m^{6})\ .$ (7)
Here we have defined the coordinate-space logarithm
$L=L_{m}=\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4},\ L_{\mu}=\ln\frac{z^{2}\mu^{2}e^{2\gamma_{E}}}{4}\ .$ (8)
If we consider large Euclidean momentum expansion, then the result reads
$S(p^{2})=\frac{1}{p^{2}}-\frac{m^{2}}{(p^{2})^{2}}+\frac{m^{4}}{(p^{2})^{3}}+\frac{g^{2}}{32\pi^{2}(p^{2})^{3}}l^{2}+{\cal O}\left(\frac{1}{(p^{2})^{4}}\right)\ ,$ (9)
$l=l_{m}=\ln\frac{p^{2}}{m^{2}}\ ,l_{\mu}=\ln\frac{p^{2}}{\mu^{2}}\ ,$ (10)
where we also defined the momentum-space logarithms. In the rest of this section we show how the OPE calculation correctly reproduces the above and allows their generalization with high-order corrections to operator condensates.
Before moving to the OPE calculation, it is time to determine the $g^{2}$ corrections to the operators. Here we only need the corrections to $\phi^{2}$, $\phi(i\partial)^{2}\phi$ and $\phi(i\partial)^{4}\phi$. The diagrams are shown in Fig. 2. To calculate these diagrams, one needs only the following three-loop vacuum diagram
$I=\frac{1}{(2\pi)^{6}}\int\frac{d^{2}kd^{2}ld^{2}q}{(k^{2}+m^{2})(l^{2}+m^{2})(q^{2}+m^{2})((k+l+q)^{2}+m^{2})}$
$=\frac{1}{(4\pi)^{3}m^{2}}\int_{0}^{\infty}\frac{d\alpha_{1}d\alpha_{2}d\alpha_{3}d\alpha_{4}}{\alpha_{1}\alpha_{2}(\alpha_{3}+\alpha_{4})+\alpha_{3}\alpha_{4}(\alpha_{1}+\alpha_{2})}e^{-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}}$
$=\frac{7\zeta_{3}}{(4\pi)^{3}m^{2}}\ .$ (11)
To obtain the last result, one splits the denominator using Mellin-Barnes. Thus, one has
$\langle\phi^{2}\rangle_{g^{2}}=-\frac{g^{2}}{6}\frac{1}{4}\partial_{m^{2}}I=\frac{g^{2}}{6}\frac{7\zeta_{3}}{(4\pi)^{3}4(m^{2})^{2}}\ ,$ (12)
$\langle\phi(i\partial)^{2}\phi\rangle_{g^{2}}=\frac{g^{2}}{6}I-m^{2}\langle\phi^{2}\rangle_{g^{2}}=\frac{g^{2}}{6}\frac{3}{4}I\ .$ (13)
The above are just the two terms for the $\langle\phi^{2}\rangle_{g^{2}}$ and $\langle\phi\partial^{2}\phi\rangle_{g^{2}}$ contributions. One also needs the divergent condensate $\langle\phi\partial^{4}\phi\rangle_{g^{2}}$ regularized in DR. One has
$\langle\phi\partial^{4}\phi\rangle_{g^{2}}=\frac{g^{2}}{6}I_{0}^{3}-2m^{2}\langle\phi(i\partial)^{2}\phi\rangle_{g^{2}}-m^{4}\langle\phi^{2}\rangle_{g^{2}}$
$=\frac{g^{2}}{6}I_{0}^{3}-\frac{g^{2}}{6}\frac{7\zeta_{3}}{(4\pi)^{3}}\bigg{(}\frac{1}{4}+\frac{3}{2}\bigg{)}+{\cal O}(\epsilon)\ ,$ (14)
where $I_{0}=\frac{\mu_{0}^{2-d}}{(2\pi)^{D}}\int\frac{d^{D}k}{k^{2}+m^{2}}$ is just the one-loop tadpole. Clearly, the first two operators are already finite at $g^{2}$ order, while the $\phi(\partial)^{4}\phi$ contain an $\frac{1}{\epsilon^{3}}$ UV singularity that we will see cancel exactly with the IR singularities for the coordinate-space coefficient functions at the power $z^{4}$.
Figure 1: The diagrams in the full theory for two-point function up to ${\cal O}(g^{2})$. The solid lines are massive propagators with mass $m^{2}$.
Figure 2: The tree-level condensates for $\phi^{2}$, $\phi(i\partial)^{2}\phi$, $\phi\partial^{4}\phi$ and $\phi^{4}$ (upper). The order ${\cal O}(g^{2})$ condensates for $\phi^{2},\phi(i\partial)^{2}\phi$ and $\phi\partial^{4}\phi$ (lower). Operator insertions are denoted as crossed circles.
Figure 3: The coefficient functions for various operators at the order ${\cal O}(z^{0})$ and ${\cal O}(z^{2})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “external soft fields” from condensates attached to the coefficient functions. They will be amputated in the coefficient function. The tree-level coefficient $1$ for $\phi^{2}$ and $-\frac{1}{2d}$ for $\phi(i\partial)^{2}\phi$ are omitted in the figure, but are crucial for coordinates space expansion.
2.1 Leading power ($z^{0}$) and next-to-leading power ($z^{2}$)
At leading power, in the momentum space there is only a single term $\frac{1}{p^{2}}$ [^1] Notice that the momentum-space two point function in $D$-dimension in our convention has an overall factor $\mu_{0}^{2-D}$. This overall factor will not participate in the momentum-space renormalization and will be omitted in this work. In the coordinate space it participates in the momentum integral and can not be neglected.. In coordinate space, on the other hand, there is already a log
$S^{(0)}(z^{2}m^{2})=-\frac{1}{4\pi}\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}+\frac{7\zeta_{3}}{24(4\pi)^{3}}\frac{g^{2}}{m^{4}}\ .$ (15)
Can one explain the logarithm here? Clearly, the condensate $\langle\phi^{2}\rangle_{0}$ is required. One has the bare coefficient function
$C_{\rm I}^{(0)}(z^{2}\mu^{2},d)=\mu^{2-d}_{0}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{e^{ip\cdot z}}{p^{2}}=\frac{1}{(4\pi)^{\frac{d}{2}}}\Gamma\left(\frac{d}{2}-1\right)\left(\frac{z^{2}\mu_{0}^{2}}{4}\right)^{1-\frac{d}{2}}\ ,$ (16)
and the tree-level bare condensate
$\langle\phi^{2}\rangle_{0}=I_{0}=\mu_{0}^{2-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{1}{p^{2}+m^{2}}=\left(\frac{\mu_{0}^{2}}{m^{2}}\right)^{1-\frac{d}{2}}\pi^{-\frac{d}{2}}\Gamma\left(1-\frac{d}{2}\right)\ .$ (17)
Clearly, the bare OPE in the $d\rightarrow 2$ limit reproduces the full logarithm
$C_{\rm I}^{(0)}(z^{2}\mu^{2},d)+\langle\phi^{2}\rangle_{0}\bigg{|}_{d\rightarrow 2}\rightarrow-\frac{1}{4\pi}\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}\ .$ (18)
The above suggests to define the minimally subtracted condensate with $d=2-2\epsilon$, as well as renormalized coefficient functions that are separately finite as $d\rightarrow 2$:
$\langle\phi^{2}\rangle_{0}=\langle\phi^{2}\rangle_{r}+\frac{1}{4\pi\epsilon}\ ,$ (19)
$C_{{\rm I},r}^{(0)}(z^{2}\mu^{2})=C_{\rm I}^{(0)}(z^{2}\mu^{2},d)+\frac{1}{4\pi\epsilon}\bigg{|}_{d\rightarrow 2}\rightarrow-\frac{1}{4\pi}\ln\frac{z^{2}\mu^{2}e^{2\gamma_{E}}}{4}=-\frac{1}{4\pi}L_{\mu}\ .$ (20)
Clearly, the $\frac{1}{4\pi\epsilon}$ pole of the bare operator exactly cancels the IR divergence of the bare coefficient function, and it can also be regarded as the IR subtraction or IR counter-term for the coefficient function. Moreover, the above allows the generalization to include high-order corrections to the operator condensates:
$\langle\phi^{2}\rangle_{r}=\frac{1}{4\pi}\ln\frac{\mu^{2}}{m^{2}}+\langle:\phi^{2}:\rangle\ ,$ (21)
$S^{(0)}\left(z^{2}m^{2},\frac{g}{m^{2}}\right)=C^{(0)}_{{\rm I},r}(z^{2}\mu^{2})+\langle\phi^{2}\rangle_{r}=-\frac{1}{4\pi}\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}+\langle:\phi^{2}:\rangle\ .$ (22)
Thus, the $\mu^{2}$ determines the separation point between short distance and large distance fluctuations in the $\overline{\rm MS}$ scheme. The $\langle:\phi^{2}:\rangle$ start to contribute at $g^{2}$ and is given in Eq. (12). It exactly reproduces the $\zeta_{3}$ term at $z^{0}$ in the first line of Eq. (2). Higher order corrections will not contribute any new logarithms at the leading power, but only modifies this condensate. Notice that the bare coefficient function completely determines the $\frac{1}{4\pi\epsilon}$ for the operator renormalization.
We then consider the nexet-to-leading power. At this order, the expansion is still simple
$S^{(2)}(z^{2}m^{2})=\frac{z^{2}m^{2}}{4}\times\frac{2-L}{4\pi}-\frac{7\zeta_{3}}{8(4\pi)^{3}}\frac{g^{2}}{m^{2}}\frac{z^{2}}{4}\ .$ (23)
There are two condensates $\phi^{2}$ and $\phi(i\partial)^{2}\phi$ in the OPE. The bare coefficient functions follow from figure 3
$C_{\rm I}^{(2)}=-m_{0}^{2}\mu_{0}^{2-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}}=-\frac{m_{0}^{2}z^{2}}{4(4\pi)^{\frac{d}{2}}}\left(\frac{z^{2}\mu_{0}^{2}}{4}\right)^{1-\frac{d}{2}}\Gamma\left(\frac{d}{2}-2\right)\ ,$ (24)
$C_{\phi^{2}}^{(2)}=-\frac{g}{2}\mu_{0}^{2-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}}=-\frac{gz^{2}}{8(4\pi)^{\frac{d}{2}}}\left(\frac{z^{2}\mu_{0}^{2}}{4}\right)^{1-\frac{d}{2}}\Gamma\left(\frac{d}{2}-2\right)\ ,$ (25)
$C_{\phi(i\partial)^{2}\phi}^{(2)}=-\frac{z^{2}}{2d}\ .$ (26)
In the momentum space, one has
$S^{(2)}(p^{2})(p^{2})^{2}=-m_{0}^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{0}=-m_{0}^{2}-\frac{g}{8\pi\epsilon}-\frac{g}{2}\langle\phi^{2}\rangle_{r}\ ,$ (27)
where we have used the renormalization constant Eq. (19). By requiring its finiteness, one determines the renormalization of the bare mass
$m_{0}^{2}=m_{r}^{2}-\frac{g}{8\pi\epsilon}\ .$ (28)
Here we have introduced short distance renormalized mass $m_{r}^{2}$ that is $\epsilon$ dependent, but is finite in the $\epsilon\rightarrow 0$ limit. It can be regarded as the bare mass subtracted only the short-distance contribution to the mass rernomalization, but without including any effects at the IR scale $m$. Due to this, it is the most natural mass that appears in the short distance calculation and in the renormalized coefficient functions.
Here, we have seen that by requiring the IR finiteness of renormalized coefficient functions automatically determines the mass renormlaization. This also resolves the “puzzle” concerning the consistency between the DR rule that throws away scaleless integrals and the fact that the $\overline{\rm MS}$ renormalization is short distance in nature: the mass renormalization in $\phi_{2}^{4}$ comes purely from the one-loop tadpole which becomes scaleless in the massless theory. How this renormalization can be determined without computing any massive integrals? Clearly, the approach here is the way: by requiring IR finiteness of the leading-power coefficient function $C_{\rm I,r}^{(0)}$ uniquely determines the operator renormalization for $\langle\phi^{2}\rangle_{0}$, the latter then determines the mass renormalization by requiring IR finiteness of the next-to-leading power coefficient function. In fact, this pattern will persist to all powers: requiring IR finiteness of the coefficient functions automatically determines the UV poles of operators.
Here we move back to the coordinate space. To determine the renormalization of $\phi\partial^{2}\phi$, one can write
$\langle\phi(i\partial)^{2}\phi\rangle_{0}=a_{1}m_{r}^{2}+a_{2}g\langle\phi^{2}\rangle_{r}+\langle\phi(i\partial)^{2}\phi\rangle_{r}\ ,$ (29)
with $a_{1},a_{2}$ are $\frac{1}{\epsilon}$ poles. Notice that in this paper, we call purely $\frac{1}{\epsilon}$ coefficients such as $a_{1}$, $a_{2}$ that express bare operators in terms of renormalized ones as UV poles of the operators. They are simultaneously IR counter-terms for the coefficient functions. Now, requiring finiteness of renormalized coefficient functions
$C_{{\rm I},r}^{(2)}=C_{\rm I}^{(2)}+\frac{1}{4\pi\epsilon}C_{\phi^{2}}^{(2)}-\frac{z^{2}}{2d}m_{r}^{2}a_{1}\rightarrow{\cal O}(1)\ ,$ (30)
$C_{\phi^{2},r}^{(2)}=C_{\phi^{2}}^{(2)}-\frac{z^{2}}{2d}ga_{2}\rightarrow{\cal O}(1)\ .$ (31)
This uniquely determines the operator poles
$a_{1}=-\frac{1}{4\pi\epsilon}\ ,a_{2}=-\frac{1}{8\pi\epsilon}\ .$ (32)
Thus, we obtains
$C_{{\rm I},r}^{(2)}=\frac{2-L_{\mu}}{4\pi}\frac{z^{2}m_{r}^{2}}{4}\ ,$ (33)
$C_{\phi^{2},r}^{(2)}=\frac{2-L_{\mu}}{4\pi}\frac{z^{2}g}{8}\ ,$ (34)
$C_{\phi(i\partial)^{2}\phi,\ r}^{(2)}=-\frac{z^{2}}{4}\ .$ (35)
Alternatively, one has for the “‘tree-level” condensate
$\langle\phi(i\partial)^{2}\phi\rangle_{0}=-m^{2}I_{0}=-\frac{m_{r}^{2}}{4\pi\epsilon}-\frac{g}{8\pi\epsilon}I_{r}-I_{r}\left(m_{r}^{2}+\frac{g}{2}I_{r}\right)\ .$ (36)
The above is consistent with $m_{0}^{2}=m^{2}-\frac{g}{2}I_{0}$ and $I_{0}=I_{r}+\frac{1}{4\pi\epsilon}$. Thus, the renormalization constants can also be read from the above equation and agrees with the one determined from the IR-UV cancellation requirement. It is easy to show that the above coefficient functions combined with the tree-level condensates exactly reproduces the full expansion at $z^{2}$ and ${\cal O}(g^{0})$. Furthermore, from the above the RGE for the renormalized operator condensate $\langle\phi\partial^{2}\phi\rangle_{r}$ and the general solution is
$\frac{d}{d\ln\mu^{2}}\langle\phi(i\partial)^{2}\phi\rangle_{r}=-\frac{m_{r}^{2}}{4\pi}-\frac{g}{8\pi}\langle\phi^{2}\rangle_{r}\ ,$ (37)
$\langle\phi(i\partial)^{2}\phi\rangle_{r}=-I_{r}\left(m_{r}^{2}+\frac{g}{2}I_{r}\right)-\frac{g}{8\pi}\ln\frac{\mu^{2}}{m^{2}}\langle:\phi^{2}:\rangle+{\cal R}\langle\phi(i\partial)^{2}\phi\rangle\ .$ (38)
The general expansion at the power $z^{2}$ is therefore after setting $\mu=m$
$S^{(2)}\left(z^{2}m^{2},\frac{g}{m^{2}}\right)=\frac{2-L}{4\pi}\frac{z^{2}}{4}\left(m^{2}+\frac{g}{2}\langle:\phi^{2}:\rangle\right)-\frac{z^{2}}{4}{\cal R}\langle\phi(i\partial)^{2}\phi\rangle\ .$ (39)
Again, the high-order contribution only modifies the two condensates, but no new form of $z$ dependency will be generated. The order $g^{2}$ value of ${\cal R}\langle\phi\partial^{2}\phi\rangle$ is given in Eq. (13) and exactly reproduces the $\zeta_{3}$ term at the power $z^{2}$ after combining with the coefficient function $-\frac{z^{2}}{4}$.
2.2 The OPE at ${\cal O}(z^{4})$
In this subsection we check the expansion at the power $z^{4}$ or $\frac{1}{(p^{2})^{3}}$ with OPE calculations. The relevant diagrams for the coefficient functions and operators are shown in Fig. 4 and Fig. 2.
2.2.1 Momentum space OPE
Figure 4: The coefficient functions for various operators at the order ${\cal O}(z^{4})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “amputated external soft fields” to keep track the operator content of the condensates. The tree-level coefficient $-\frac{3z^{4}}{24d(d+2)}$ for $\phi\partial^{4}\phi$ are omitted in the figure, but crucial for coordinate space expansion.
To check the expansion against the OPE, one needs the coefficient functions. In the momentum space, one needs the following functions
$C_{0}\left(\frac{p^{2}}{\mu^{2}},d\right)=\mu_{0}^{4-2d}\int\frac{d^{d}kd^{d}l}{(2\pi)^{2d}}\frac{1}{k^{2}l^{2}(p-k-l)^{2}}$
$=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{2-d}\frac{8^{1-d}\pi^{\frac{3}{2}-d}\Gamma\left(\frac{d}{2}-1\right)^{2}}{\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{3d}{2}-3\right)\sin\pi d}\ ,$ (40)
$C_{1}\left(\frac{p^{2}}{\mu^{2}},d\right)=\mu_{0}^{2-d}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}(p-k)^{2}}$
$=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{1-\frac{d}{2}}\frac{2^{3-2d}\pi^{\frac{1}{2}-\frac{d}{2}}\Gamma\left(2-\frac{d}{2}\right)\Gamma\left(\frac{d}{2}-1\right)}{\Gamma\left(\frac{d-1}{2}\right)}\ ,$ (41)
$I_{0}\left(\frac{\mu^{2}}{m^{2}},d\right)=\left(\frac{\mu_{0}^{2}}{m^{2}}\right)^{1-\frac{d}{2}}\pi^{-\frac{d}{2}}\Gamma\left(1-\frac{d}{2}\right)\equiv\langle\phi^{2}\rangle_{0}\ ,$ (42)
$\langle\phi^{4}\rangle_{0}=3I_{0}^{2}\ .$ (43)
Notice that $C_{0}$ and $C_{1}$ are just the massless self-energy diagrams with zero and two external field insertions without symmetry factors, see Fig. 4. One can show that
$C_{0}\left(\frac{p^{2}}{\mu^{2}},d\right)+3C_{1}\left(\frac{p^{2}}{\mu^{2}},d\right)I_{0}\left(\frac{\mu^{2}}{m^{2}},d\right)+\frac{3}{p^{2}}I_{0}^{2}\left(\frac{\mu^{2}}{m^{2}},d\right)\bigg{|}_{d\rightarrow 2}\rightarrow\frac{3}{p^{2}(4\pi)^{2}}\ln^{2}\frac{p^{2}}{m^{2}}\ ,$ (44)
which is exactly the expansion of the massive sunrise self-energy $\Sigma(p^{2})$ defined in Eq. (6). Clearly, this can be regarded as follows: $\Sigma(p^{2})$ at the power $\frac{1}{p^{2}}$ contains the following soft sub-regions: there are three ways for one of the three propagators in the sunrise to become soft and form the quadratic condensate $\langle\phi^{2}\rangle_{0}$. There are also three ways for two of them to form the quartic condensate $\langle\phi^{4}\rangle_{0}$. Combining with the region where all three propagators are hard, one indeed reproduces the expansion for $\Sigma(p^{2})$ at $\frac{1}{p^{2}}$.
We would like to proceed further to obtain the OPE for the two-point function. For this one must notice that the $\langle\phi^{4}\rangle_{0}$ insertion for the self-energy is not all of the $\langle\phi^{4}\rangle_{0}=3I_{0}^{2}$ contribution to the two-point function. The reason is that there is also a “reducible” term that should cancel with the $\frac{1}{\epsilon^{2}}$ for double $m_{0}^{2}$ insertion. Indeed, if one compute the last diagram in Fig. (4) as a correlation function between $\phi(-p)\phi(p)$ with four amputated zero-momentum scalar fields, then one can see the coefficient function for $\phi_{0}^{4}$ insertion should be $\frac{g^{2}}{4p^{2}}$. The contribution to the self-energy is $\frac{g^{2}}{6p^{2}}\times 3I_{0}^{2}$, while there is another $\frac{1}{12p^{2}}\times 3I_{0}^{2}$ that should cancel with mass insertions. Similarly, for the $\langle\phi^{2}\rangle_{0}$ insertion one should also include the $m_{0}^{2}\times g$ order diagram shown in Fig. 4. Therefore, the bare OPE for the two-point function reads
$S(p^{2})(p^{2})^{2}=\left(\frac{g^{2}}{6}C_{0}+\frac{m_{0}^{4}}{p^{2}}\right)+\left(\frac{g^{2}}{2}C_{1}+\frac{gm_{0}^{2}}{p^{2}}\right)\langle\phi^{2}\rangle_{0}+\left(\frac{g^{2}}{6p^{2}}+\frac{g^{2}}{12p^{2}}\right)\langle\phi^{4}\rangle_{0}\ .$ (45)
Then, one introduces the renormalized operators in the $\overline{MS}$ scheme as
$\langle\phi^{2}\rangle_{0}=\langle\phi^{2}\rangle_{r}+\frac{1}{4\pi\epsilon}\ ,$ (46)
$\langle\phi^{4}\rangle_{0}=\langle\phi^{4}\rangle_{r}+\frac{3}{2\pi\epsilon}\langle\phi^{2}\rangle_{r}+\frac{3}{(4\pi)^{2}\epsilon^{2}}\ .$ (47)
Notice that $\langle\phi^{4}\rangle_{r}=3\langle\phi^{2}\rangle_{r}^{2}$ at tree-level, despite possible $\frac{1}{\epsilon}\times\epsilon$ effects. The point is that $\langle\phi^{2}\rangle_{r}$ and $\langle\phi^{4}\rangle_{r}$ are finite in the $\epsilon\rightarrow 0$ limit, but in the $\epsilon\rightarrow 0$ limit they can still contain ${\cal O}(\epsilon)$ terms.
Given the above, by collecting coefficients for renormalized operators, one has
$S(p^{2})(p^{2})^{2}=C_{{\rm I},r}^{(4)}(p)+C_{\phi^{2},r}^{(4)}(p)\langle\phi^{2}\rangle_{r}+C_{\phi^{4},r}^{(4)}(p)\langle\phi^{4}\rangle_{r}\ ,$ (48)
$C_{{\rm I},r}^{(4)}(p)=\frac{g^{2}}{6}C_{0}+\frac{g^{2}}{8\pi\epsilon}C_{1}+\frac{g^{2}}{2p^{2}}\frac{1}{(4\pi)^{2}\epsilon^{2}}+\frac{m_{r}^{4}}{p^{2}}\ ,$ (49)
$C_{\phi^{2},r}^{(4)}(p)=\frac{g^{2}}{2}C_{1}+\frac{g^{2}}{2p^{2}}\frac{1}{2\pi\epsilon}+\frac{gm_{r}^{2}}{p^{2}}\ ,$ (50)
$C_{\phi^{4},r}^{(4)}(p)=\frac{g^{2}}{4p^{2}}\ .$ (51)
In the $\epsilon\rightarrow 0$ limit, all of the above are finite. In fact, requiring that the $\epsilon\rightarrow 0$ limit exist uniquely determines all the $\frac{1}{\epsilon}$ poles in the mixing, once the tree-level mixing constant $Z_{10}^{(0)}=\frac{1}{4\pi\epsilon}$ for $\phi^{2}$ is determined from the coordinate space expansion at $z^{0}$ order, and the one-loop level mixing $m_{0}^{2}=m_{r}^{2}-\frac{g}{8\pi\epsilon}$ is fixed from the $z^{2}$ order consideration. In particular
$C_{{\rm I},r}^{(4)}(p)=\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}+\frac{m_{r}^{4}}{p^{2}}\ ,$ (52)
$C_{\phi^{2},r}^{(4)}(p)=\frac{g^{2}}{4\pi p^{2}}\ln\frac{p^{2}}{\mu^{2}}+\frac{gm_{r}^{2}}{p^{2}}\ .$ (53)
Notice that all the $m_{0}^{2}$ insertions in $C_{\rm I}$ again combine with the operator UV poles for $\phi^{2n}$ inserted along the same propagator to form the renormalized mass $m_{r}^{2}$. Anticipating that, in the coefficient function calculation, one can instead use $m_{r}^{2}$ and neglect many operator UV poles corresponding to “reducible” insertions of $(\phi^{2})^{n}$ along the same propagator. This is particularly convenient in the large $N$ calculation, since in such cases different contractions between the $(\phi^{2})^{n}$ usually have different large $N$ power-counting.
Given the OPE, to reproduce the full result, notice in the current case one has $m_{r}^{2}=m^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{r}=m^{2}-\frac{g}{8\pi}\ln\frac{\mu^{2}}{m^{2}}+{\cal O}(\epsilon)$, and the tree-level operator condensates in the $\epsilon\rightarrow 0$ limit read
$\langle\phi^{2}\rangle_{r}=\frac{1}{4\pi}\ln\frac{\mu^{2}}{m^{2}}\ ,$ (54)
$\langle\phi^{4}\rangle_{r}=\frac{3}{16\pi^{2}}\ln^{2}\frac{\mu^{2}}{m^{2}}\ .$ (55)
Combining all above, one has the $\mu$-independent result $S_{2}(p^{2})(p^{2})^{2}=\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{m^{2}}+\frac{m^{4}}{p^{2}}=C_{{\rm I},r}^{(4)}(\mu^{2}=m^{2})$. This again reveals an important fact: if only tree-level condensates are required, then one can set $\mu=m$ to remove them and keep only the coefficient function for the identity operator. This is because at tree-level, $\langle\phi^{2n}\rangle$ always factorize into $\frac{(2n)!}{(n!)2^{n}}\langle\phi^{2}\rangle^{n}$, therefore after renormalization they always vanish at $\mu=m$. This is also true if some of the $\phi^{2}$s are replaced by $\phi(\partial^{2k}\phi)$, but only self-contractions are present, and only the “tadpole-type” high-order corrections to these operators are allowed in terms of the $\frac{1}{k^{2}+m_{r}^{2}}$ propagators. Later in Sec. (3) we will use this observation.
Here we should stress that if one only keeps the naive perturbative result $\frac{g^{2}}{6}C_{0}$ in $C_{{\rm I},r}$ without introducing additive counter-terms, then not only the $\frac{1}{\epsilon}$ poles no longer cancel, but the double log term in the finite part would also have the wrong coefficient $\frac{g^{2}}{16\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}$.
$\frac{g^{2}}{6}C_{0}=\frac{g^{2}}{32\pi^{2}p^{2}\epsilon^{2}}-\frac{g^{2}}{16\pi^{2}p^{2}\epsilon}\ln\frac{p^{2}}{\mu^{2}}+\frac{g^{2}}{16\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}-\frac{g^{2}}{32\pi^{2}p^{2}}\zeta_{2}+{\cal O}(\epsilon)\ .$ (56)
In the “additive counter-term” $\frac{g^{2}}{8\pi\epsilon}C_{1}$, the $\frac{1}{\epsilon}\times\epsilon$ terms will generate another $-\frac{g^{2}}{32\pi^{2}p^{2}}\ln\frac{p^{2}}{\mu^{2}}$ :
$\frac{g^{2}}{8\pi\epsilon}C_{1}=-\frac{g^{2}}{16\pi^{2}p^{2}\epsilon^{2}}+\frac{g^{2}}{16\pi^{2}p^{2}\epsilon}\ln\frac{p^{2}}{\mu^{2}}-\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}+\frac{g^{2}}{32\pi^{2}p^{2}}\zeta_{2}+{\cal O}(\epsilon)\ .$ (57)
It is clear now that all the $\frac{1}{\epsilon}$ poles cancel, the $\frac{1}{\epsilon^{2}}$ poles add to cancel the remaining $\frac{g^{2}}{2p^{2}}\frac{1}{(4\pi)^{2}\epsilon^{2}}$. The $\zeta_{2}$ term all cancel and the double-log terms combine to reproduce the full result for $C_{{\rm I},r}$. Therefore, in the DR calculation, it is important to keep the $d$ dependency of the bare coefficient functions to the end, and the $\epsilon\times\frac{1}{\epsilon}$ effects can be crucial.
Beyond the order $g^{2}$, the condensates will be modified, but the full expansion will be scale invariant using the RGE
$\frac{d}{d\ln\mu^{2}}\phi_{r}^{2}=\frac{1}{4\pi}\ ,$ (58)
$\frac{d}{d\ln\mu^{2}}\phi_{r}^{4}=\frac{3}{2\pi}\phi_{r}^{2}\ .$ (59)
Thus, one has
$\langle\phi^{2}\rangle_{r}=\frac{1}{4\pi}\ln\frac{\mu^{2}}{m^{2}}+\langle:\phi^{2}:\rangle\ ,$ (60)
$\langle\phi^{4}\rangle_{r}=\frac{3}{16\pi^{2}}\ln^{2}\frac{\mu^{2}}{m^{2}}+\frac{3}{2\pi}\ln\frac{\mu^{2}}{m^{2}}\langle:\phi^{2}:\rangle+\langle:\phi^{4}:\rangle\ .$ (61)
It is not difficult to show that here $\langle:\phi^{2}:\rangle$ and $\langle:\phi^{4}:\rangle$ are just the vacuum condensates for $\mu$-independent Wick-polynomials. Given the $\mu$-independent condensates, the expansion will be modified as
$S(p^{2})(p^{2})^{2}=\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{m^{2}}+\frac{m^{4}}{p^{2}}+\bigg{(}\frac{gm^{2}}{p^{2}}+\frac{g^{2}}{4\pi p^{2}}\ln\frac{p^{2}}{m^{2}}\bigg{)}\langle:\phi^{2}:\rangle+\frac{g^{2}}{4p^{2}}\langle:\phi^{4}:\rangle\ .$ (62)
Here we note that the above is the general form of the large $p^{2}$ expansion at the next-to-next-leading order in $\frac{1}{p^{2}}$ for a generic $\phi_{2}^{4}$, where $\langle:\phi^{2}:\rangle$ and $\langle:\phi^{4}:\rangle$ can be non-perturbative. At small $g$, condensates for these two Wick-polynomials are of orders ${\cal O}(\frac{g^{2}}{m^{4}})$ and ${\cal O}\left(\frac{g}{m^{2}}\right)$. The double logarithmic dependency is always purely determined perturbatively.
2.2.2 Coordinate space OPE
We now move back to the coordinate space. The coefficient functions can be Fourier transformed into the $z$ space. Here one needs the following functions
$C_{0}(z^{2}\mu^{2},d)$ $=\mu_{0}^{6-3d}\int\frac{d^{d}pd^{d}kd^{d}l}{(2\pi)^{3d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}k^{2}l^{2}(p-k-l)^{2}}$
$=z^{4}(z^{2}\mu_{0}^{2})^{\frac{6-3d}{2}}\frac{2^{-d-5}\pi^{-\frac{3}{2}(d-1)}\Gamma\left(\frac{d}{2}-1\right)^{2}}{(9d^{2}-54d+80)\Gamma(5-d)\Gamma\left(\frac{d-1}{2}\right)\sin\pi d}\ ,$ (63)
$C_{1}(z^{2}\mu^{2},d)$ $=\mu_{0}^{4-2d}\int\frac{d^{d}pd^{d}k}{(2\pi)^{2d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}k^{2}(p-k)^{2}}$
$=-z^{4}(z^{2}\mu_{0}^{2})^{2-d}\frac{2^{-d-5}\pi^{\frac{3}{2}-d}\Gamma(d-4)}{\Gamma\left(4-\frac{d}{2}\right)\Gamma\left(\frac{d-1}{2}\right)\sin\left(\frac{\pi d}{2}\right)}\ ,$ (64)
$C_{2}(z^{2}\mu^{2},d)$ $=\mu_{0}^{2-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{e^{ip\cdot z}}{(p^{2})^{3}}$
$=z^{4}(z^{2}\mu_{0}^{2})^{1-\frac{d}{2}}\frac{\pi^{-\frac{d}{2}}}{128}\ \Gamma\left(\frac{d}{2}-3\right)\ ,$ (65)
$C_{3}(z^{2},d)$ $=\frac{3z^{4}}{24d(d+2)}\ .$ (66)
Notice that they are already for the two-point function rather than the self-energy. Given the above, one has
$C_{0}(z^{2}\mu^{2},d)+3C_{1}(z^{2}\mu^{2},d)I_{0}\left(\frac{\mu^{2}}{m^{2}},d\right)+3C_{2}(z^{2}\mu^{2},d)I_{0}^{2}\left(\frac{\mu^{2}}{m^{2}},d\right)+C_{3}(z^{2},d)I_{0}^{3}\left(\frac{\mu^{2}}{m^{2}},d\right)$
$\bigg{|}_{d\rightarrow 2}\rightarrow\frac{z^{4}}{16(4\pi)^{3}}\bigg{(}-\frac{L_{\mu}^{3}}{4}+\frac{9L_{\mu}^{2}}{4}-\frac{69L_{\mu}}{8}+\frac{27}{2}\bigg{)}\ ,$ (67)
which is exactly the “hard” contribution in Eq. (2), including the constant $\frac{27}{2}$. Notice that the last term is due to the UV divergent $I_{0}^{3}$ contribution in the $\langle\phi\partial^{4}\phi\rangle_{g^{2}}$ as shown in Eq. (14) and the $\frac{3z^{4}}{24d(d+2)}$ is the coefficient obtained by Taylor expanding
$\langle\phi\partial^{i}\partial^{j}\partial^{k}\partial^{l}\phi\rangle=\frac{\delta^{ij}\delta^{kl}+\delta^{ik}\delta^{jl}+\delta^{il}\delta^{jk}}{d(d+2)}\langle\phi\partial^{4}\phi\rangle\ .$ (68)
Namely, one should regard $\frac{3z^{4}}{24d(d+2)}$ as the coefficient for the scalar operator $\langle\phi\partial^{4}\phi\rangle$.
We now rewrite the OPE in coordinate space in terms of renormalized operators. For this purpose one needs the renormalization of $\phi\partial^{4}\phi$, which can be read from below with $I_{r}=I_{0}-\frac{1}{4\pi\epsilon}=\langle\phi^{2}\rangle_{r}$
$m^{4}I_{0}+\frac{g^{2}I_{0}^{3}}{6}=$ $\frac{m_{r}^{4}}{4\pi\epsilon}+\frac{g^{2}}{384\pi^{3}\epsilon^{3}}+\bigg{(}\frac{g^{2}}{32\pi^{2}\epsilon^{2}}+\frac{gm_{r}^{2}}{4\pi\epsilon}\bigg{)}\langle\phi^{2}\rangle_{r}+\frac{g^{2}}{16\pi\epsilon}\langle\phi^{4}\rangle_{r}$
$+\frac{g^{2}I_{r}^{3}}{6}+I_{r}\left(m_{r}^{2}+\frac{g}{2}I_{r}\right)^{2}+{\cal O}(\epsilon)\ ,$ (69)
which defines the UV poles and renormalized operators. Defining the operators
$O_{0}=1\ ,O_{1}=\phi^{2},O_{2}=\phi^{4},O_{3}=\phi\partial^{4}\phi\ ,$ (70)
then for zero momentum transfer, the RGE are two-loop exact and read
$\frac{\partial O_{i}}{\partial\ln\mu^{2}}=\gamma_{ij}O_{j}\ ,$ (71)
$\gamma_{0i}=0\ ,\ \gamma_{10}=\frac{1}{4\pi}\ ,\gamma_{21}=\frac{3}{2\pi}\ ,\ \gamma_{30}=\frac{m_{r}^{4}}{4\pi}\ ,\ \gamma_{31}=\frac{gm_{r}^{2}}{4\pi}\ ,\ \gamma_{32}=\frac{g^{2}}{16\pi}\ .$ (72)
Given the above, one has the renormalized coefficient functions
$C_{{\rm I},r}^{(4)}=\frac{g^{2}}{6}C_{0}+\frac{g^{2}}{8\pi\epsilon}C_{1}+\frac{g^{2}}{4\pi^{2}\epsilon^{2}}\frac{C_{2}}{2}+m_{r}^{4}C_{2}+\left(\frac{g^{2}}{384\pi^{3}\epsilon^{3}}+\frac{m_{r}^{4}}{4\pi\epsilon}\right)C_{3}\ ,$ (73)
$C_{\phi^{2},r}^{(4)}=\frac{g^{2}}{2}C_{1}+\frac{g^{2}}{2\pi\epsilon}\frac{C_{2}}{2}+gm_{r}^{2}C_{2}+\bigg{(}\frac{g^{2}}{32\pi^{2}\epsilon^{2}}+\frac{gm_{r}^{2}}{4\pi\epsilon}\bigg{)}C_{3}\ ,$ (74)
$C_{\phi^{4},r}^{(4)}=\frac{g^{2}}{4}C_{2}+\frac{g^{2}}{16\pi\epsilon}C_{3}\ ,$ (75)
$C_{\phi(i\partial)^{4}\phi,r}^{(4)}=C_{3}\ .$ (76)
In the $\epsilon\rightarrow 0$ limit, all the above are finite. The explicit results read
$C_{{\rm I},r}^{(4)}=\frac{1}{(4\pi)^{3}}\bigg{(}-\frac{L_{\mu}^{3}}{24}+\frac{3L_{\mu}^{2}}{8}-\frac{23L_{\mu}}{16}+\frac{9}{4}\bigg{)}\left(\frac{z^{2}g}{4}\right)^{2}+\frac{(3-L_{\mu})}{16\pi}\left(\frac{z^{2}m_{r}^{2}}{4}\right)^{2}\ ,$ (77)
$C_{\phi^{2},r}^{(4)}=\frac{1}{(4\pi)^{2}}\bigg{(}\frac{L_{\mu}^{2}}{8}-\frac{3L_{\mu}}{4}+\frac{23}{16}\bigg{)}\left(\frac{z^{2}g}{4}\right)^{2}+\frac{(3-L_{\mu})g}{16\pi}\left(\frac{z^{2}m_{r}}{4}\right)^{2}\ ,$ (78)
$C_{\phi^{4},r}^{(4)}=\frac{1}{4}\frac{(3-L_{\mu})}{16\pi}\left(\frac{z^{2}g}{4}\right)^{2}\ ,$ (79)
$C_{\phi(i\partial)^{4}\phi,r}^{(4)}=\frac{z^{4}}{64}\ .$ (80)
Here in the $L_{\mu}$, $m$ is replaced by the $\overline{MS}$ scheme $\mu$. In fact, by requiring the cancellation of $\frac{1}{\epsilon}$ poles, the three series of UV poles for $O_{3}\rightarrow O_{i}$ mixing
$Z_{30}=\frac{g^{2}}{384\pi^{3}\epsilon^{3}}+\frac{m_{r}^{4}}{4\pi\epsilon}\ ,$ (81)
$Z_{31}=\frac{g^{2}}{32\pi^{2}\epsilon^{2}}+\frac{gm_{r}^{2}}{4\pi\epsilon}\ ,$ (82)
$Z_{32}=\frac{g^{2}}{16\pi\epsilon}\ ,$ (83)
can all be uniquely determined, once the renormalization of $m_{0}$ and $\langle\phi^{2}\rangle$ are fixed by $z^{0}$ and $z^{2}$ power considerations.
It is easy to see that the OPE coefficients above satisfies the RGE $\frac{d}{d\ln\mu}C_{i}=-C_{j}\gamma_{ji}$ and the full expansion is RGE invariant. In particular, to order $g^{2}$, the expansion of $S(z^{2}m^{2},g^{2})$ at the order $z^{4}m^{4}$ read with $\mu=m$
$S(z^{2}m^{2},g^{2})=$ $\frac{1}{(4\pi)^{3}}\bigg{(}-\frac{L^{3}}{24}+\frac{3L^{2}}{8}-\frac{23L}{16}+\frac{9}{4}\bigg{)}\left(\frac{z^{2}g}{4}\right)^{2}+\frac{(3-L)}{16\pi}\left(\frac{z^{2}m^{2}}{4}\right)^{2}$
$-\frac{z^{4}}{64}\times\frac{g^{2}}{6}\times\frac{49\zeta_{3}}{4(4\pi)^{3}}\ ,$ (84)
where the last term is exactly due to the finite part of the non-trivial condensate of $\phi\partial^{4}\phi$ at order $g^{2}$ in Eq. (14). Beyond the $g^{2}$ order, all contributions are included in the three non-trivial condensates. In particular, one has
$\langle\phi\partial^{4}\phi\rangle_{r}=$ $\frac{g^{2}I_{r}^{3}}{6}+I_{r}\left(m_{r}^{2}+\frac{g}{2}I_{r}\right)^{2}+\left(\frac{gm^{2}}{4\pi}\ln\frac{\mu^{2}}{m^{2}}+\frac{g^{2}}{32\pi^{2}}\ln^{2}\frac{\mu^{2}}{m^{2}}\right)\langle:\phi^{2}:\rangle$
$+\frac{g^{2}}{16\pi}\ln\frac{\mu^{2}}{m^{2}}\langle:\phi^{4}:\rangle+{\cal R}\langle\phi\partial^{4}\phi\rangle\ ,$ (85)
where $I_{r}=\frac{1}{4\pi}\ln\frac{\mu^{2}}{m^{2}}+{\cal O}(\epsilon)$ and ${\cal R}(\phi\partial^{4}\phi)=-\frac{g^{2}}{6}\frac{49\zeta_{3}}{4(4\pi)^{3}}+{\cal O}(g^{3})$. It is also interesting to notice that by Fourier transforming the coordinate space expansion term by term, one obtains the momentum-space expansion. On the other hand, from the momentum-space expansion, one can not Fourier-transforming back, unless additional subtraction terms at $\vec{p}=0$ are included to make the expansion tempered.
3 Higher order behavior of OPE expansion in a large-$N$ quartic $O(N)$ model: factorial enhancements and their cancellation
Given the example in the previous section, we can see that the renormalized coefficient functions for various operators can be calculated in massless fixed-point perturbation theory with IR subtractions. One of the natural questions is, what is the high-order behavior of the coefficient functions, when regarded as an asymptotic expansion in $z^{2}g$?
For example, let’s consider the coefficient function $C_{\rm I}$ for the identity operator that arises from $\phi(z)\times\phi(0)$. The $C_{\rm I}$ has an expansion
$C_{\rm I}=\sum_{n=0}^{\infty}z^{2n}g^{n}\sum_{k=0}^{n}\left(\frac{m_{r}^{2}}{g}\right)^{k}C_{n,k}(L_{\mu})\ ,$ (86)
where $C_{n,k}$ depends only on $L_{\mu}$. Here we focus on the terms $C_{n,k}$. These terms are nothing but the IR renormalized two-point functions in the massless perturbation theory at the order $n-k$ with $k$ mass insertions. Heuristically, one may argue that the total numbers of Feynman diagrams grow as $(n)!$, while being the $n$-th power in the small $z^{2}$ expansion introduces another factor of $\frac{1}{(n!)^{2}}$. Thus $C_{n,k}$ may still be suppressed by $\frac{1}{n!}$, suggesting an infinite convergence radius for the power expansion. However, the fact that $C_{n,k}$ requires IR-renormalization, may introduce additional factorial growth, in a way similar to the famous IR-renormalon problem ’t Hooft (1979); Parisi (1979, 1978) for the logarithmic asymptotics in marginal theory.
To investigate this issue, one may again use the large-$N$ expansion trick. Consider the following $O(N)$ generalization of the scalar $\phi^{4}$ theory in 2D  [^2] In Mariño and Reis (2020), this model has been studied in the $O(N)$-broken false vacuum and a renormalon has been observed in the coupling-constant expansion of the free-energy. In our work, we are interested in the OPE expansion in the physical unbroken phase, and it remains to see if there are connections between the renormalon found in Mariño and Reis (2020) and the factorial enhancements observed in our work.
$S=\mu_{0}^{D-2}\int d^{D}x\bigg{(}\frac{1}{2}\partial_{i}\Phi^{a}\partial_{i}\Phi^{a}+\frac{m^{2}+\delta m^{2}}{2}\Phi^{a}\Phi^{a}+\frac{g}{8N}:(\Phi^{a}\Phi^{a})::(\Phi^{a}\Phi^{a}):\bigg{)}\ .$ (87)
Here $a=1,2,3,...N$ with summation convention implied, and we have use again the bare mass $m_{0}^{2}=m^{2}+\delta m^{2}-\frac{g}{2}I_{0}$. Here, the Wick-ordering is performed with respect to the free Gaussian field with mass $m$. Notice that $\delta m^{2}={\cal O}\left(\frac{g}{N}\right)$ which we chose to cancel the one-loop self-energy correction at $\frac{1}{N}$. Equivalently, one may write in terms of the overall Wick-ordering $:(\Phi^{a}\Phi^{a})^{2}:$. Also notice that the normalization for the quartic term is different from that of the $N=1$. Unlike the $N=1$ case, for $N\geq 3$ it is normally conjectured in a way similar to the lattice $O(N)$ model Polyakov (1975) that there is no phase transition as one increases the coupling constant. To simplify notation, from now on we use
$\Phi^{2}\equiv\sum_{a=1}^{N}\Phi^{a}\Phi^{a}\ ,\ \Phi(i\partial)^{n}\Phi\equiv\sum_{a=1}^{N}\Phi^{a}(i\partial)^{n}\Phi^{a}\ ,$ (88)
to denote $O(N)$-contracted quadratic operators.
Here we consider the coefficient function for the identity operator in the OPE exapsion in the short distance limit $z^{2}g\ll 1,z^{2}m^{2}\ll 1$, for again the two point function $\langle\Phi^{a}(z)\Phi^{b}(0)\rangle\rightarrow\delta^{ab}C_{\rm I}(z)+....$. The IR structure is irrelevant for this limit. Clearly, one has the large $N$ expansion $C_{\rm I}=C_{\rm I,L}+\frac{1}{N}C_{\rm I,NL}+...$ To calculate $C_{\rm I,{\rm NL}}$, notice that we only need to consider a massless bubble chain and subtract out the IR sub-divergences to obtain the IR renormalized coefficient functions. To further amplify the factorial enhancements, and to make contact with previous investigations, in the following we perform the calculation in the momentum space expansion, but will comment on the features of the coordinate space expansion.
3.1 Coefficient functions from IR renormalized bubble chain
To be more precise, we consider the $\frac{1}{N}$ order two-point function in momentum-space
$S(p^{2})(p^{2}+m^{2})^{2}=\frac{g}{N}\int\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(p-k)^{2}+m^{2}}\frac{gF(k^{2})}{1+gF(k^{2})}+\frac{g}{2N}\langle:\Phi^{2}:\rangle\ ,$ (89)
$F(k^{2})=\frac{1}{2}\int\frac{d^{2}l}{(2\pi)^{2}}\frac{1}{(l^{2}+m^{2})((k-l)^{2}+m^{2})}=\frac{1}{4\pi k^{2}}\frac{1}{A(k^{2})}\ln\frac{A(k^{2})+1}{A(k^{2})-1}\ ,$ (90)
$A(k^{2})=\sqrt{1+\frac{4m^{2}}{k^{2}}}\ .$ (91)
Clearly, all contributions are absolutely convergent. The factor $\frac{gF}{1+gF}$ clearly has the interpretation of a massive chain of bubbles starting with bubble number $n=1$. The $n=0$ contribution is logarithmically divergent and subtracted by $\delta m^{2}$. We would like to examine the OPE expansion of the above. First notice that the last term in Eq. (89) is due to the $\frac{1}{N}$ order tadpole contribution and contains no factorial behavior at all, so we focus on the the first term. See Fig. 5 for a depiction of the massive bubble chain diagram with $n=5$.
Figure 5: The massive bubble-chain diagram with $n=5$ bubbles. Solid lines are massive scalar propagators. Notice in the figure we have separated the quartic interaction vertices slightly apart just to show the contractions among the $O(N)$ indices, no point-spiting-type regularization has been implied. At the order $\frac{1}{N}$ for the two-point function $S(p^{2})$, the $O(N)$ indices only contract within each bubbles.
By investigating the regions, the operators beside the identify operator that we will need are the following:
  1. 1.
    Tree-level operators $\prod_{l}\frac{1}{N}\Phi(i\partial)^{2k_{l}}\Phi$. At the order $\frac{1}{N}$, these operators are responsible for the IR sub-contributions within the bubble chain when $k^{2}$ itself is in the hard region, as well as the IR contribution on the $(p-k)$ line when $p-k$ becomes soft (this means $k$ is hard). Clearly, the “tree-level” insertions within the bubble-chain are mutually independent and can be performed separately within the individual bubbles. Indeed, when $k$ is hard, only one line within each bubble can become soft, and the Taylor expansion for that soft momentum do not affect other bubbles. To perform the insertion responsible for the small $p-k$ region, one must Taylor expand the momentum flowing into the bubble-chain, but the condensates formed by making $p-k$ soft are still the tree-level condensates and do not correlate with the tree-level condensates formed within the bubble chain. At the external legs and any hard internal lines, there can also be an arbitrary number of tree-level $\frac{1}{N}\Phi^{2}$ insertions that always combine together with the $m_{r}^{2}$ insertion. All these tree-level condensates can be removed by simply setting $\mu^{2}=m^{2}$.
  2. 2.
[…]
    $\frac{1}{(p-k)^{2}}=\sum_{n=0}^{\infty}\frac{k^{2n}}{(p^{2})^{n+1}}\frac{\Gamma(1-\epsilon)\Gamma(\epsilon+n+1)}{\Gamma(1+\epsilon)\Gamma(-\epsilon+n+1)}\ ,$ (92)
    -- $\frac{m^{2}}{(p-k)^{4}}=-\frac{m^{2}}{p^{2}}\sum_{n=0}^{\infty}\frac{k^{2n}}{(p^{2})^{n+1}}\frac{(n+1)\Gamma(1-\epsilon)\Gamma(\epsilon+n+2)}{\Gamma(2+\epsilon)\Gamma(-\epsilon+n+1)}\ ,$ (93)
    and with more mass insertions one can generalize without difficulty. In Appendix. B, we explain how such angular averages can be performed. These coefficient functions are clearly absent from any factorial growth. See Fig. 6 for a depiction of these operator contributions.
Figure 6: The contribution of the operator $O_{n}$ due to the soft $k^{2}$ region, in the presence of zero (upper) and one (lower) mass-insertions after expanding the $\frac{1}{(p-k)^{2}+m^{2}}$ at small $k$. The dashed line represents massless scalar propagator, black dot denotes the mass insertion and the gray-blob denotes the operator condensate $\langle O_{n}\rangle$. As the case of $\phi_{2}^{4}$, double-lines represent “amputated soft external fields” to keep track the operators associated to the coefficient functions.
Given the above and using the fact that all tree-level operators can be removed simply by setting $\mu=m$, the expansion at $\mu=m$ takes a very simple form
$S_{\rm NL}(p^{2})-\frac{g}{2(p^{2}+m^{2})^{2}}\langle:\Phi^{2}:\rangle=C_{{\rm I},{\rm NL}}(p^{2})+\sum_{n=1}^{\infty}\frac{\langle O_{n}\rangle}{(p^{2})^{n}}\frac{\,{}_{2}F_{1}\left(n,n;1;-\frac{m^{2}}{p^{2}}\right)}{(p^{2}+m^{2})^{2}}\ ,$ (94)
where $\langle O_{n}\rangle$ denotes the renormalized operator condensates at $\mu^{2}=m^{2}$. Notice that the disconnected contributions to $O_{1}$ also vanish at $\mu^{2}=m^{2}$ to the desired $\frac{1}{N}$ order. To obtain the last term in Eq. (94), one needs to expand the $\frac{1}{(p-k)^{2}+m^{2}}$ to the order $k^{2n-2}$ and perform the angular average, see Appendix. B for more details. For other values of $\mu$, there will be many tree-level operators as discussed above, but the coefficient function $C_{\rm I;{\rm NL}}$ remains essentially the same form with $m_{r}^{2}$ replacing $m^{2}$ in the pre-coefficients, while $\mu^{2}$ replacing $m^{2}$ in the logarithms. For coefficient functions of ${\cal O}_{n}$, one simply replacing $m$ with $m_{r}$. From the above, it is clear that the coefficient functions for $O_{n}$ contain no factorial growths at all. Thus, to search for factorial growths, it is sufficient to start from the coefficient function for the identity operator. We can write the $\frac{1}{N}$ order coefficient function of the identity operator for $S_{\rm NL}(p^{2})$ in the general form as
$C_{{\rm I},{\rm NL}}(p^{2})=\frac{1}{(p^{2}+m_{r}^{2})^{2}}\sum_{n=1}^{\infty}C^{(n)}_{\rm I,NL}(p^{2})\ ,$ (95)
$C^{(n)}_{\rm I,NL}(p^{2})=\frac{g^{n+1}}{(p^{2})^{n}}\sum_{k=0}^{n-1}\left(\frac{m_{r}^{2}}{g}\right)^{k}C_{n,k}(l_{\mu})\ ,$ (96)
where $C_{n,k}$ is a polynomial on $l_{\mu}=\ln\frac{p^{2}}{\mu^{2}}$. Notice $C_{n,n}=0$. Since merely mass insertion will not cause the factorial growth, as can be checked from simple calculations in lower-order diagrams, we start from $C_{n,0}$.
3.1.1 Calculation of $C_{n,0}$: factorial growth from $\frac{1}{\epsilon}\times\epsilon$
Figure 7: The massless bubble chain diagram contributing to $C_{n,0}$ with $n=5$ without mass insertions (upper) and the IR subtractions (lower) excluding the one for $O_{n}$. Since the contraction of $O(N)$ indices is clear, we no longer separate the quartic interactions in the figure. The IR subtractions on multiple bubbles factorize and can be performed separately in each bubble. Similarly, the IR counter-term for the $p-k$ line also factorizes.
Given the above, let’s consider the calculation detail. For individual bubble chain, one needs the building block that already appeared earlier in Eq. (41)
$\frac{1}{2}C_{1}=\frac{\mu_{0}^{2-d}}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}(p-k)^{2}}=\frac{1}{p^{2}}\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}f(\epsilon)\times\frac{-1}{4\pi\epsilon}\ ,$ (97)
where
$f(\epsilon)=\frac{\sqrt{\pi}2^{2\epsilon}e^{\gamma_{E}\epsilon}\Gamma(1-\epsilon)\Gamma(1+\epsilon)}{\Gamma\left(\frac{1}{2}-\epsilon\right)}=1-\frac{\pi^{2}}{12}\epsilon^{2}+{\cal O}(\epsilon^{3})\ .$ (98)
Given the above, the bare coefficient function (in the below we use hat to denote bare or un-subtracted coefficient functions) corresponding to the $n$-bubble chain diagram (depicted in the upper of Fig. 7 with $n=5$) reads
$\frac{g^{n+1}}{(p^{2})^{n}}\hat{C}_{n,0}=(-1)^{n+1}(-1)^{n}\frac{g^{n+1}}{(4\pi)^{n}}\frac{f^{n}(\epsilon)}{\epsilon^{n}}\times\mu^{2\epsilon}\times\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{\mu^{2}}{k^{2}}\right)^{n\epsilon}\frac{1}{(k^{2})^{n}}\frac{1}{(p-k)^{2}}$
[…]
$=-\frac{g^{n+1}}{(4\pi)^{n+1}}\frac{f^{n}(\epsilon)}{\epsilon^{n}}e^{\epsilon\gamma_{E}}\times\frac{1}{(p^{2})^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\Gamma(-\epsilon)\Gamma(1-n-s)\Gamma(n+s)}{\Gamma(1-n-\epsilon-s)\Gamma(n+s-\epsilon)}\ ,$ (99)
where we have introduced the factor $s=(n+1)\epsilon$. The bared coefficient function thus can be expressed as
$\hat{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\frac{G(n,\epsilon,s)}{\epsilon^{n+2}(n+1)}\ ,$ (100)
where
$G(n,\epsilon,s)=f^{\frac{s}{\epsilon}-1}(\epsilon)e^{\epsilon\gamma_{E}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\epsilon\Gamma(-\epsilon)s\Gamma(1-n-s)\Gamma(n+s)}{\Gamma(1-n-\epsilon-s)\Gamma(n+s-\epsilon)}\ .$ (101)
The crucial property of the above is that it is analytic at $s=0$ and $\epsilon=0$. Now, the IR subtraction for each of the bubble can be performed by adding the operator UV poles $\frac{1}{4\pi\epsilon}$ in Eq. (19) for $\frac{1}{N}\Phi^{2}$ insertions replacing one of the two propagators in $C_{1}$ (clearly there are two ways, thus the subtraction is $2\times\frac{1}{2}\times\frac{1}{4\pi\epsilon}\times\frac{1}{p^{2}}$, where $\frac{1}{p^{2}}$ is another propagator Taylor-expanded to the zeroth order):
$\frac{1}{2}C_{1}\rightarrow\frac{1}{p^{2}}\times\frac{-1}{4\pi\epsilon}\left(\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}f(\epsilon)-1\right)\ .$ (102)
Therefore, the bubble-chain-partly-renormalized diagram can be obtained by replacing $s\rightarrow(n+1-k)\epsilon$ with $k$ ranging from 00 to $n$ and summing over $k$ using the binomial theorem:
$R_{b}\hat{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}\ .$ (103)
Notice we used the notation $R_{b}F$ to denote the bubble-chain-partly-renormalized quantity $F$, for which the IR sub-divergences within the bubble chain (not including the entire bubble-chain) have been subtracted. See Fig. 7 for a depiction of the bubble-chain diagram and IR subtractions for $C_{n,0}$ with $n=5$. Now, one needs the standard summation formulas Palanques-Mestre and Pascual (1984); Beneke and Braun (1994, 1995), as well as the new one (this is actually related to the second Stirling number $S(n+1,n)$)
$\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k+1)^{n+1}=\frac{(n+2)!}{2}\ .$ (104)
To show this, one writes
$\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k+x)^{n+1}$
$=x(n+1)\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n}+\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n+1}$
$=x(n+1)n!+n!\binom{n+1}{2}=\frac{n!(n+1)(n+2x)}{2}\ .$ (105)
This is due to the fact that $\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n+1}$ is nothing but $n!S(n+1,n)$ where $S(n+1,n)$ is the Stirling number of the second kind counting the partition of $n+1$ objects into $n$ unlabeled sets. Given this summation formula, one obtains up to all constant terms,
$R_{b}\hat{C}_{n,0}$
$=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{(-1)^{n}}{(n+1)\epsilon^{n+2}}G(n,\epsilon,0)+\frac{1}{(n+1)\epsilon}\frac{d^{n+1}}{ds^{n+1}}G(n,\epsilon,s)+\frac{1}{2}\frac{d^{n+2}}{ds^{n+2}}G(n,\epsilon,s)\bigg{)}_{s=0}\ .$ (106)
The term that causes the trouble is the second one. We need
$G(n,\epsilon,s)$
$=-s\left(\frac{\mu^{2}}{p^{2}}\right)^{s}+\frac{\epsilon}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{(}-24s\gamma_{E}+\pi^{2}s^{2}-12s\psi(1-n-s)-12s\psi(n+s)\bigg{)}+{\cal O}(\epsilon^{2})\ .$ (107)
Now, the first term will contribute to a $\ln p^{2}$-dependent $\frac{1}{\epsilon}$ pole,
$-\frac{1}{(n+1)\epsilon}\times\frac{d^{n+1}}{ds^{n+1}}s\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{|}_{s=0}=-\frac{1}{\epsilon}\ln^{n}\frac{\mu^{2}}{p^{2}}\ .$ (108)
This term must be canceled by subtracting the IR contribution from the $p-k$ propagator. Since here the IR counter term is again the UV pole $\frac{1}{4\pi\epsilon}$ for the $\frac{1}{N}\Phi^{2}$ operator, one simply needs to multiply this UV pole with the IR-subtracted bubble chain at the momentum $k=p$. Therefore, one has
$T_{p-k}R_{b}\hat{C}_{n,0}=-\frac{1}{4\pi\epsilon}\times\frac{1}{(4\pi)^{n}\epsilon^{n}}\left(f(\epsilon)\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}-1\right)^{n}\ .$ (109)
Here, $T_{p-k}F$ denotes the quantity obtained from $F$ such that the $p-k$ line have been replaced by the appropriate IR counter-term. Expanding the bracket using the binomial theorem, after using the summation formulas one reaches
$T_{p-k}R_{b}\hat{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{1}{\epsilon}\frac{d^{n}}{ds^{n}}H(\epsilon,s)\bigg{|}_{s=0}+\frac{n+2}{2}\frac{d^{n+1}H(\epsilon,s)}{ds^{n+1}}\bigg{|}_{s=0}\bigg{)},$
$H(\epsilon,s)=f^{\frac{s}{\epsilon}-1}(\epsilon)\bigg{(}\frac{\mu^{2}}{p^{2}}\bigg{)}^{s-\epsilon}\ .$ (110)
To further simplify, one needs the small $\epsilon$ epansion
$H(\epsilon,s)=\left(\frac{\mu^{2}}{p^{2}}\right)^{s}-\frac{\epsilon}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\left(12\ln\left(\frac{\mu^{2}}{p^{2}}\right)+\pi^{2}s\right)+{\cal O}(\epsilon^{2})\ .$ (111)
The $p$-dependent divergent terms are now
$\frac{1}{\epsilon}\frac{d^{n}}{ds^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{|}_{s=0}=\frac{1}{\epsilon}\ln^{n}\frac{\mu^{2}}{p^{2}}\ ,$ (112)
which indeed cancel with that for $R_{b}C_{n,0}$. Given the above, the partly renormalized $C_{n,0}$, including all IR subtractions except the one for the operator $O_{n}$ reads
$R_{p}\hat{C}_{n,0}=R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (113)
It remains to add the operator poles for $O_{n}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n-2}\Phi^{2}$. Clearly, since all the high-order derivatives in $s$ are finite in $R_{p}C_{n,0}$, the operator poles must cancel the remaining $p$-independent divergences in the first term of $R_{b}C_{n,0}$ after the equality in Eq. (3.1.1). Therefore, one writes fully renormalized $C_{n,0}$ as
$C_{n,0}=R_{p}\hat{C}_{n,0}-\frac{1}{(4\pi)^{n+1}}Z_{n}\left(\frac{1}{\epsilon}\right)\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\ ,$ (114)
where $-\frac{Z_{n}g^{n+1}}{(4\pi)^{n+1}}$ denotes operator UV poles for the $O_{n}\rightarrow{\rm I}$ mixing that are at the order $g^{n+1}$. See Fig. 8 for a depiction of the IR subtractions corresponding to UV poles of $O_{n}$. Clearly, since the dimension of $O_{n}$ is $4+2n-2=2(n+1)$, this term has the highest power in $g$ in the UV poles of $O_{n}$. We further separate the high-order derivatives in $s$ and constant terms in $C_{n,0}$:
$C_{n,0}=C_{n,0}^{A}+C_{n,0}^{B}\ ,$ (115)
where $C_{n,0}^{B}$ contains all the high-order derivatives in $s$ and is finite (its final form will be given in Eq. (127)), while
$C_{n,0}^{A}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{(-1)^{n}}{(n+1)\epsilon^{n+2}}G(n,\epsilon,0)+Z_{n}\left(\frac{1}{\epsilon}\right)\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{)}\ .$ (116)
Now, for $n\in Z_{\geq 1}$ one can simplifies the above as
$C_{n,0}^{A}=-\frac{1}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{(}Z_{n}+\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ .$ (117)
Clearly, the above uniquely determines the $Z$ as the pole part of
$Z_{n}=-\text{poles of}\ \bigg{(}\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ ,$ (118)
and the renormalized constant parts are therefore
$C_{n,0}^{A}=-\frac{1}{(4\pi)^{n+1}}\text{finite of}\ \bigg{(}\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ .$ (119)
Clearly, this part is free from any factorial growth since the function $\frac{\sin(\pi\epsilon)\Gamma(-2\epsilon)}{\Gamma^{2}(1-\epsilon)}$ is analytic in a neighborhood of $\epsilon=0$, thus the Taylor expansion coefficients can only growth exponentially.
Figure 8: The IR counter-terms due to UV poles of the operator $O_{n}$, in the presence of zero (upper) and one (lower) mass-insertions after expanding the $\frac{1}{(p-k)^{2}+m^{2}}$ at small $k$. The crossed circle denotes the $g^{n+1}$ order operator UV poles of the $O_{n}\rightarrow{\rm I}$ mixing. These contributions should be added back to $C_{n,0}$ and $C_{n+1,1}$ respectively.
Now, one collect all the terms that are due to high order derivatives in $s$. One has
$C_{n,0}^{B}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{1}{n+1}\frac{d^{n+1}}{ds^{n+1}}{\cal H}_{1}(s)+\frac{d^{n}}{ds^{n}}{\cal H}_{2}(s)+\frac{1}{2}\frac{d^{n+2}}{ds^{n+2}}{\cal H}_{3}(s)+\frac{n+2}{2}\frac{d^{n+1}}{ds^{n+1}}{\cal H}_{4}(s)\bigg{)}_{s=0}\ .$ (120)
Here
${\cal H}_{1}(s)=\frac{1}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{(}-24s\gamma_{E}+\pi^{2}s^{2}-12s\psi(1-n-s)-12s\psi(n+s)\bigg{)}\ ,$ (121)
${\cal H}_{2}(s)=-\frac{1}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\left(12\ln\left(\frac{\mu^{2}}{p^{2}}\right)+\pi^{2}s\right)\ ,$ (122)
${\cal H}_{3}(s)=-s\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\ ,$ (123)
${\cal H}_{4}(s)=\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\ .$ (124)
Notice that the first two terms arise due to $\frac{1}{\epsilon}\times\epsilon$ effects. For $n=1$, one obtains $C_{1,0}^{A}=0$, while
$C_{1,0}^{B}=C_{1,0}=\frac{1}{(4\pi)^{2}}\frac{3}{2}\ln^{2}\frac{p^{2}}{\mu^{2}}\ ,$ (125)
which agrees exactly with the previous calculation. For $n=2$, one has
$C_{2,0}=-\frac{1}{(4\pi)^{2+1}}\bigg{(}\frac{4l_{\mu}^{3}}{3}-2l_{\mu}^{2}-4l_{\mu}+\frac{10\zeta_{3}}{3}-4\bigg{)}\ ,$ (126)
where $l_{\mu}=\ln\frac{p^{2}}{\mu^{2}}$. It is clear from now that $C_{n,0}^{A}$ and ${\cal H}_{2}$ to ${\cal H}_{4}$ terms will not contribute to factorial growth at all. However, the ${\cal H}_{1}$ contribution contains the very dangerous $s\psi(1-n-s)+s\psi(n+s)$ term, which can contribute to factorial growth.
To determine the leading factorial behavior, one notices that the contribution $C_{n,0}^{B}$ can be simplified after many cancellations:
$C_{n,0}^{B}(l_{\mu})=\frac{n!}{(4\pi)^{n+1}}\oint\frac{ds}{s^{n+1}}e^{-sl_{\mu}}\bigg{(}2\gamma_{E}-l_{\mu}+\psi(1-n-s)+\psi(n+s)\bigg{)}\ .$ (127)
Given the above, it is clear that the leading non-Borel summable large $n$ asymptotic can be extracted by shifting the contours outside and pick up the negative residues at $s=1$
$C_{n,0}\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{\mu^{2}}{p^{2}}\bigg{)}\times\left(1+\frac{\mu^{2}}{2^{n+1}p^{2}}+..\right)\ .$ (128)
Notice that here we introduced the following notation to distinguish alternating and non-alternating asymptotics in $n$: for alternating asymptotics of a quantity $F(n)$ we write $F(n)|_{n\rightarrow\infty,(-1)^{n}}$, while for non-alternating one we write $F(n)|_{n\rightarrow\infty,+1}$. It is interesting to notice that the leading factorial growth is again one more power higher. Here there are two possibilities. The first possibility is that they cancel within the same coefficient function for the identity operator but with mass-insertions, if $\mu^{2}$ is identified as the mass. We will rule out this possibility after calculating the diagrams with mass-insertions. The second possibility is that they cancel with the operator contributions for $O_{n+1}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n}\Phi^{2}$ and $O_{n+2}$…., but with the same bubble-chain number, because this diagram contains power-divergences for these operators. This would be consistent with the interpretation that the non-alternating asymptotics in Eq. (128) can be regarded as an IR renormalon.
It is also interesting to notice that there are also alternating factorial growths
$C_{n,0}\bigg{|}_{n\rightarrow\infty,(-1)^{n}}=\frac{(-1)^{n}}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{p^{2}}{\mu^{2}}\bigg{)}\times\left(1+\frac{p^{2}}{2^{n+1}\mu^{2}}+...\right)\ .$ (129)
Later we will show that they cancel with factorial growths for operators $O_{n-1},O_{n-2}...$ but with the same numbers of bubbles.
3.1.2 Mass insertion: $C_{n,1}$
To rule out the possibility that the renormalon above could cancel with the diagrams with mass insertion, let’s consider $C_{n+1,1}$, namely coefficient functions with one mass insertions. There are three possibilities. Either the insertion is within the bubbles, or it is on the $p-k$ line, or it is on the external lines. Since the last type of contributions are trivial and has already been factorized out in our definition, here we only focus on the first two types of mass-insertions. In order to label the mass insertions, here we introduced the new notation: $C_{n;l,k}$ denotes the contribution to $C_{n,l+k}$ where $l$ mass insertions are on the $p-k$ line, while $k$ mass insertions are within the bubble chain. One clearly has $C_{n,k}=\sum_{l=0}^{k}C_{n;l,k-l}$.
Figure 9: The massless bubble chain diagram contributing to $C_{n+1;0,1}$, with one mass insertion within the bubbles for $n=5$ (upper) and the related IR subtractions (lower). Notice that the IR subtraction for the mass-inserted propagator is due to the $a_{1}$ in Eq. (32).
We first consider the mass insertion within the bubble chain, see Fig. 9 for a depiction. Notice that the mass-insertion diagram can be placed in $n$ locations in the bubble-chain with the same result, one can simply place it at the end of the bubble chain with a total number of $n$ bubbles. The bubble with one mass insertion reads
$C_{m}(k^{2})=-m_{r}^{2}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\int\frac{d^{D}l}{(2\pi)^{D}}\frac{1}{(k-l)^{2}}\frac{1}{(l^{2})^{2}}$
[…]
$=\frac{m_{r}^{2}}{k^{4}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\frac{(1+2\epsilon)f(\epsilon)}{2\pi\epsilon}\ ,$ (130)
with the same $f(\epsilon)$ as in Eq. (98). Here we need to perform the IR subtraction to $C_{m}$. Clearly, the $\frac{1}{(k-l)^{2}}$ propagator can still be re-normalized using the $\frac{1}{4\pi\epsilon}$ UV pole for $\frac{1}{N}\Phi^{2}$, while the $\frac{-m_{r}^{2}}{(l^{2})^{2}}$ propagator has to be renormalized by the UV pole for the $\frac{1}{N}\Phi(i\partial)^{2}\Phi$ operator insertion replacing the $\frac{1}{(l^{2})^{2}}$ line, using the Eq. (92) for the small-$l$ expansion rule of $\frac{1}{(k-l)^{2}}$ and Eq. (32) for the UV pole $a_{1}m_{r}^{2}$, which is exactly proportional to $m_{r}^{2}$. Therefore, the IR subtraction for the $C_{m}$ can be performed as
$\frac{m_{r}^{2}}{k^{4}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\frac{(1+2\epsilon)f(\epsilon)}{2\pi\epsilon}+\frac{-m_{r}^{2}}{k^{4}}\times\frac{1}{4\pi\epsilon}+\frac{1}{k^{4}}\frac{1+\epsilon}{1-\epsilon}\times a_{1}m_{r}^{2}$
$=\frac{2m_{r}^{2}}{k^{4}}\frac{f(\epsilon)}{4\pi\epsilon}\left(f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\right)+\frac{2m_{r}^{2}}{k^{4}}\frac{1}{4\pi}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}2f(\epsilon)-\frac{2m_{r}^{2}}{k^{4}}\frac{1}{4\pi}g(\epsilon)\ ,$ (131)
where $g(\epsilon)$ is defined as
$g(\epsilon)=-\bigg{(}\frac{1}{\epsilon}-\frac{1}{2\epsilon}-\frac{1+\epsilon}{2\epsilon(1-\epsilon)}\bigg{)}\equiv\frac{1}{(1-\epsilon)}\ .$ (132)
It is again regular at $\epsilon=0$. Given the above, one can combine with the rest $n-1$ bubbles without the mass-insertion to obtain the bubble-chain-partly-renormalized diagram
$R_{b}\hat{C}_{n+1;0,1}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
$\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{1}{\epsilon^{n}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n}\frac{1}{(p-k)^{2}}$
$+\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{2f(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}$
$-\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{g(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}\ .$ (133)
We can further split the second term to obtain
$\frac{f(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}$
$=\frac{1}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n}\frac{1}{(p-k)^{2}}$
$+\frac{1}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}\ .$ (134)
The bubble-chain-partly-renormalized $C_{n+1;0,1}$ therefore reads
$R_{b}\hat{C}_{n+1;0,1}=$ $\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{2G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+1}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n-1}(-1)^{k}\binom{n-1}{k}\frac{(2-g(\epsilon))G_{1}(n,\epsilon,(n-k)\epsilon)}{\epsilon^{n+1}(n-k)}\ .$ (135)
Here the $G_{1}$ function is defined as
$G_{1}(n,\epsilon,s)=e^{\gamma_{E}\epsilon}f(\epsilon)^{\frac{s}{\epsilon}-1}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\epsilon\Gamma(-\epsilon)s\Gamma(-n-s)\Gamma(n+s+1)}{\Gamma(-\epsilon-n-s)\Gamma(-\epsilon+n+s+1)}\equiv G(n+1,\epsilon,s)\ .$ (136)
The $\ln p$-dependent $\frac{1}{\epsilon}$ terms are generated only from the first and the last two terms in Eq. (3.1.2), which are
$\frac{2n}{(4\pi)^{n+1}}\bigg{(}-\frac{1}{n+1}\frac{d^{n+1}}{ds^{n+1}}s\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{|}_{s=0}\bigg{)}$
$+\frac{2n}{(4\pi)^{n+1}}\bigg{(}-\frac{1}{n}\frac{d^{n}}{ds^{n}}s\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{|}_{s=0}\bigg{)}=\frac{2n}{(4\pi)^{n+1}}\left(-\ln^{n}\frac{\mu^{2}}{p^{2}}-\ln^{n-1}\frac{\mu^{2}}{p^{2}}\right)\ .$ (137)
It is not hard to show that after adding the counter-term $T_{p-k}R_{b}\hat{C}_{n+1;0,1}$ for the $p-k$ line, all the $\ln p$ in the $\frac{1}{\epsilon}$ cancel, leading to $\ln p$-independent poles for the operator . This time, the operator poles determined are the $g^{n+1}m_{r}^{2}$ terms for $O_{n+1}\rightarrow{\rm I}$ (or $g^{n}m_{r}^{2}$ term for $O_{n}\rightarrow{\rm I}$). The diagram for this operator pole can still be read from the upper of Fig. 8, in which $-\frac{Z_{n}g^{n+1}}{(4\pi)^{n+1}}$ should be replaced by the $m_{r}^{2}g^{n+1}$-order poles for the $O_{n+1}$. In fact, by staying at the same $n$ but considering all the $C_{n;0,k}$ with $10\leq k\leq n-$ 1, one determines all the operator poles for the $O_{n}\rightarrow{\rm I}$ mixing. One can proceed further using the same method as in the case of $C_{n,0}$ to obtain the fully renormalized $C_{n+1;0,1}$. Since this part of calculation is very similar to that of $C_{n,0}$, here we omit further details of it. The final result implies that there are also factorial growths in $C_{n;0,1}$. In fact, the leading non-alternating factorial growth reads:
$C_{n;0,1}\bigg{|}_{n\rightarrow\infty,+1}=\frac{2n!}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}-.....\ .$ (138)
Thus, for generic values of $\frac{g}{m^{2}}$ it will not cancel the factorial growth of $C_{n,0}$.
Figure 10: The massless bubble chain diagram contributing to $C_{n+1;1,0}$, with one mass insertion one the $p-k$ line $n=5$ (upper) and the related IR subtractions (lower). Notice that the IR subtraction for the mass-inserted $p-k$ propagator is again due to the $a_{1}$ in Eq. (32). Notice that here although the IR counter-term $a_{1}$ do not correlate with IR counter-terms within the bubble chain, to perform the IR subtraction one must Taylor expand the IR subtracted bubble chain to the second order, see last term in Eq. (139).
The mass-insertion on the $\frac{1}{(p-k)^{2}}$ line can also be calculated similarly, see Fig. 10 for a depiction. This contribution is more illustrative than the mass-insertion within the bubble chain, since it requires the same operator poles as that of $O_{n}$ determined from the $C_{n,0}$ calculation, therefore it serves as a consistency check for the IR subtraction method. Here we can write directly the partly renormalized result (partly-renormalized-bubble-chain plus the IR counter-term on $p-k$)
$R_{p}\hat{C}_{n+1;1,0}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
$\frac{m_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{1}{\epsilon^{n}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n}\frac{1}{((p-k)^{2})^{2}}$
$+\frac{m_{r}^{2}g^{n+1}}{(4\pi)^{n+1}}\frac{1}{\epsilon^{n+1}}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}-1\bigg{)}^{n}\ .$ (139)
Notice that the last term is due to the subtraction of IR divergence in the small $p-k$ region using the UV pole $a_{1}$ in Eq. (32) for the $\frac{1}{N}\Phi(i\partial)^{2}\Phi$ operator. Clearly, in order to perform this subtraction, one must Taylor expand the IR subtracted bubble chain to the second order, explaining the second line of Eq. (139). To further simplify, one needs the function
$G_{2}(n,\epsilon,s)$ $=f(\epsilon)^{\frac{s}{\epsilon}-1}e^{\epsilon\gamma_{E}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\epsilon\Gamma(-1-\epsilon)s\Gamma(1-n-s)\Gamma(1+n+s)}{\Gamma(-n-\epsilon-s)\Gamma(n+s-\epsilon)}$
$=\frac{(n+s)(n+s+\epsilon)}{1+\epsilon}G(n,\epsilon,s)\ .$ (140)
The pre-factor $\frac{(n+s)(n+s+\epsilon)}{1+\epsilon}$ suggests to write
$\frac{1}{(p^{2})^{n+1}}G_{2}(n,\epsilon,s)=\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}G(n,\epsilon,s)\ .$ (141)
In terms of the above, one has
$\frac{1}{(p^{2})^{n+1}}R_{p}\hat{C}_{n+1;1,0}=$ $-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}R_{b}\hat{C}_{n,0}$
$-\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (142)
Here $R_{b}\hat{C}_{n,0}$ is given in Eq. (3.1.1) and $T_{p-k}R_{b}\hat{C}_{n,0}$ is given in Eq. (3.1.1).
To proceed, one splits the second line of Eq. (3.1.2) to obtain
$\frac{1}{(p^{2})^{n+1}}R_{p}\hat{C}_{n+1;1,0}=$ $-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\left(R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\right)$
$-\frac{2\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (143)
From this equation, it is clear that the $\frac{1}{\epsilon}$ terms from the high-order derivatives again cancel. This is consistent with the fact that after subtracting all the IR sub-divergences, one only needs $p$-independent operator poles. In fact, the required operator poles here are still the $g^{n+1}$ terms in $Z_{n}$ for $O_{n}\rightarrow{\rm I}$, but attached to the $m_{r}^{2}$ power coefficient function:
$\frac{-Z_{n}}{(4\pi)^{n+1}}\times\frac{1}{(p^{2})^{n+1}}\frac{-n\Gamma(1-\epsilon)}{\Gamma(2+\epsilon)}\frac{\Gamma(n+\epsilon+1)}{\Gamma(n-\epsilon)}$
$=\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\frac{Z_{n}}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)}{\Gamma(1+\epsilon)}\frac{\Gamma(n+\epsilon)}{\Gamma(n-\epsilon)}\ .$ (144)
See lower part of Fig. 8 for a depiction of this operator pole. Therefore, the contribution obtained by combining the operator poles and the first line of Eq. (143) in the $\epsilon\rightarrow 0$ limit, exactly leads to $C_{n,0}$ inside the derivatives
$-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\left(R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\right)$
$+\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\frac{Z_{n}}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)}{\Gamma(1+\epsilon)}\frac{\Gamma(n+\epsilon)}{\Gamma(n-\epsilon)}$
$\rightarrow-\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{C_{n,0}}{(p^{2})^{n}}\ .$ (145)
Here we have used Eq. (114) for fully renormalized $C_{n,0}$. Therefore, using the $Z_{n}$ already given in Eq. (118), we have seen that it automatically renormalizes the coefficient function with one more mass insertion on the $p-k$ line. This is a crucial consistency check for our IR subtraction. On the other hand, the $\frac{2\epsilon}{1+\epsilon}$ term in the second line of Eq. (143) combines with the $\frac{1}{\epsilon}$ term in the $T_{p-k}R_{b}\hat{C}_{n,0}$ given in Eq. (3.1.1) to form a finite contribution.
Combining all above, the fully renormalized $C_{n+1;1,0}$ reads
$\frac{1}{(p^{2})^{n+1}}C_{n+1;1,0}$
$=-\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{C_{n,0}}{(p^{2})^{n}}+\frac{1}{(4\pi)^{n+1}}\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{2}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}\ .$ (146)
Clearly, the factorial growth is from the firs term. Using the previous result Eq. (128) for $C_{n,0}$, one therefore obtains the leading non-alternating asymtotics as
$C_{n;1,0}\bigg{|}_{n\rightarrow\infty,+1}=\frac{1}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}n\times n!+....\ .$ (147)
Clearly, not only it will not cancel the contribution in $C_{n,0}$, but it has a factor-$n$ enhancement in comparison to $C_{n,0}$. The enhancement is clearly due to the fact that the mass-insertion is roughly equivalent to two derivatives. One can proceed further without difficulty and show that the leading factorial growth of $C_{n;l,0}$, namely, all the $l$ mass-insertions are on the $p-k$ line, reads
$C_{n;l,0}\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1-l}}\frac{\mu^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ .$ (148)
Later we will see that the above summed over $l$ is almost the leading large-$n$ asymptotics of $C_{\rm I,NL}^{n}$, up to an over-all multiplication constant $e^{-\frac{8\pi m_{r}^{2}}{g}}$ that is exponentially small in $g$. For now, simply notice that the maximal value of the asymptotics above is achieved around $l=\sqrt{n}$, with is roughly $n!\frac{\sqrt{\frac{1}{n}}e^{2\sqrt{n}-\frac{1}{2}}}{2\pi}$.
3.2 Cancellation of factorial growth between coefficient functions and operators
Given the discussion in the previous subsection, it is clear from now that the factorial growths in the coefficient functions can not be canceled within the coefficient functions. Instead, they are expected to cancel with the operators. How to see the factorial growth in the operators? We first consider the alternating asymptotics of $C_{n,0}$ in Eq. (129). It looks like an operator contribution from $O_{n-1}$ but with $n$-bubbles. Let’ show it is indeed the case. For this we need
$\langle O_{n-1}\rangle_{g^{n+1}}=(-1)^{n+1}\int\frac{d^{2}k}{(2\pi)^{2}}(k^{2})^{n-2}g(gF(k^{2}))^{n}\ .$ (149)
Here we focus on the large-momentum region, where one has the logarithmic behavior
$F(k^{2})\rightarrow\frac{1}{4\pi k^{2}}\ln\frac{k^{2}}{m^{2}}\ ,$ (150)
thus
$\langle O_{n-1}\rangle_{g^{n+1}}$ $\rightarrow(-1)^{n+1}\frac{g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\gg\alpha m^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{4}}\ln^{n}\frac{k^{2}}{m^{2}}$
$\rightarrow\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n+1}}\frac{n!}{m^{2}}\ .$ (151)
Interestingly, it exactly cancels the leading alternating factorial growth in Eq. (129) after setting $\mu=m$. For other $\mu$, how this can still be explained? Notice that although $O_{n-1}$ in the $n$ bubble diagram is UV finite, for other $\mu$ there are sub-divergences due to iterated insertion of tadpoles, which must be renormalized and leads to polynomials in $\ln\frac{\mu^{2}}{m_{r}^{2}}$. In particular, if expressed in $m_{r}$ and $\mu$, then there must be a contribution to $O_{n-1}$ of the form
$\langle O_{n-1}(\mu)\rangle_{n+1}=\frac{g^{n+1}}{m_{r}^{2}}\sum_{k=0}^{n}o_{k}\ln^{k}\frac{\mu^{2}}{m_{r}^{2}}\ ,$ (152)
the $k=0$ term is just the UV finite contribution calculated before, while for $k\geq 1$ one replaces $k$ bubble-chains with tadpole-type insertions. The $o_{k}$ for small $k$ contains factorial growth in $n$, but alternates between different $k$. The combined effect is that there is a factorial growth of the form $\frac{1}{m_{r}^{2}}\times\frac{m_{r}^{2}}{\mu^{2}}$ that exactly matches the $\frac{1}{\mu^{2}}$ in the Eq. (129). In conclusion, the alternating “UV renormalons” in $C_{n,0}$ have to be canceled by lower-dimension operators.
We then moves to the non-alternating factorial growths Eq. (128). They resemble operator contributions from $O_{n+1}$ but still with $n$-bubbles, thus there are UV divergences which should be renormalized according to our renormalization schemes. There is a contribution of the order $m_{r}^{2}g^{n+1}$, while the logarithms could lead to $g^{n+1}m_{r}^{2}\times\frac{\mu^{2}}{m_{r}^{2}}n!=g^{n+1}\mu^{2}n!$. The fact that $\mu^{2}$ is in the numerator indicates that this is a UV renormalon for the operator. To show this, one can use the following trick. Let’s consider the point-splitting regularization
$\frac{g^{2}}{4N}\langle\Phi^{2}(z)(i\partial)^{2n}\Phi^{2}(0)\rangle_{c}=\frac{1}{z^{2}}(C_{{\rm I},0}(L)+...+m^{2}z^{2}C_{{\rm I},1}(L)+...)+\langle O_{n+1}\rangle+....\ .$ (153)
Now, by subtracting out all the singular contributions at $\frac{1}{z^{2}}$, as well as finite contributions from the mass-insertions, one reaches the $O_{n}$. So one can investigate if these subtractions could introduce factorial growths. Since we are interested in the factorial growth that correlates with the coefficient functions for the identity operator with no mass insertions, one focus on $C_{{\rm I},0}$. In fact, one has here the bared one for the $n$ bubble-contribution in it
$\frac{1}{z^{2}}\hat{C}_{n,0}(L)=-\frac{g^{n+1}e^{\gamma_{E}\epsilon}}{(4\pi)^{n+1}}\frac{4}{z^{2}}\left(\frac{z^{2}\mu^{2}}{4}\right)^{s}\frac{f(\epsilon)^{n}\Gamma(1-s)}{\epsilon^{n}\Gamma(s-\epsilon)}\ ,$ (154)
it contains exactly a $s=1$ renormalon singularity that is none-alternating. In the large $n$ limit, the factorial growth after the IR subtraction is therefore
$\frac{1}{z^{2}}C_{n,0}(L)\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}}n!\mu^{2}\ .$ (155)
Therefore, after subtracting the $C_{{\rm I},0}(L)$ in Eq. (153), an opposite factorial growth is generated:
$\langle O_{n+1}\rangle\bigg{|}_{n\rightarrow\infty,+1}=\frac{g^{n+1}}{(4\pi)^{n+1}}n!\mu^{2}\ .$ (156)
Combining with its tree-level coefficient function, it cancels exactly the leading factorial growth in Eq. (128).
Similarly, one can show that the factorial growths for the $m_{r}^{2}$ terms in the coefficient functions are canceled by two sources. First, the $n!$ growth of $C_{n;0,1}$ from mass insertions within the bubble chains are canceled by the $O_{n}$ operator in $n-1$ bubbles. This time, the renormalon is proportional to $m_{r}^{2}$ and can be obtained in the point-splitting by subtracting the $m_{r}^{2}$ order coefficient function $C_{{\rm I},1}$ from the $\frac{1}{z^{4}}(C_{{\rm I},0}+m_{r}^{2}z^{2}C_{{\rm I},1})$. Second, the $n\times n!$ growth of $C_{n;1,0}$ from $m_{r}^{2}$ insertion on the $p-k$ line can be explained by the operator $O_{n}$ in $n-1$ bubbles but attached to the $m_{r}^{2}$ order coefficient function. Indeed, the small-$k$ expansion of $-\frac{m_{r}^{2}}{(p-k)^{4}}$ introduces the factor of $-n^{2}m_{r}^{2}$. This combined with the operator’s $(n-1)!\mu^{2}\frac{g^{n}}{(4\pi)^{n}}$ exactly cancels the factorial growth in Eq. (147). So the cancellation of leading factorial growth between coefficient functions and operators is checked up to $m_{r}^{2}$.
To conclude, the cancellation pattern is: alternating “renomalons” cancel with lower-dimensional operators, while non-alternating renormalons cancel with higher-dimensional operators. The renormalons always cancel at the same numbers of $g$, and the cancellation is always at the same values of $s$. For example, the $s=2$ subleading renormalon corresponding to the second term in Eq. (128) would cancel with the sub-leading UV renormalon of the operator $O_{n+2}$ with $n$-bubbles (and diagrams that can be obtained by deleting bubbles but adding tadpole-type insertions).
3.3 Checking the OPE calculation with Mellin-Barnes
Finally, in this subsection we show that the result obtained in the previous subsections can also be derived from Mellin-Barnes representation based on Eq. (89).
The strategy is, one first introduces $10<s_{1}<$ 1 and write
$\frac{gF(k^{2})}{1+gF(k^{2})}=\int_{c-i\infty}^{c+i\infty}\frac{ds_{1}}{2\pi i}\frac{\pi}{\sin\pi s_{1}}(gF(k^{2}))^{s1}\ .$ (157)
We then write
[…]
which should be understood as analytically continued from the $s_{1}<0$ region. At this step, one can introduces another Mellin-Barnes parameters $s$ in a way similar to those in Beneke et al. (1998); Marino and Miravitllas (2024); Liu (2024)
$A^{-s_{1}}\left(\frac{A-1}{A+1}\right)^{t}=\int_{-t<\rm Re{s}<\frac{1}{2}s_{1}}\frac{ds}{2\pi i}\frac{\,{}_{2}\tilde{F}_{1}(s_{1}-1,s+t;-s+s_{1}+t;-1)}{\sin 2\pi s\Gamma(2s-s_{1}+1)\Gamma(-s-t+1)}$
×$\times\pi^{2}\csc(\pi(s-s_{1}-t))(\sin(\pi s_{1})\csc(\pi(2s-s_{1}))-\sin(\pi(s-t))\csc(\pi(s+t)))\left(\frac{k^{2}}{m^{2}}\right)^{s}\ .$ (159)
We then write
$\frac{\pi}{\sin\pi s_{1}}\left(\frac{g}{4\pi}\right)^{s_{1}}g\int\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{s_{1}}}\left(\frac{k^{2}}{m^{2}}\right)^{s}\frac{1}{(p-k)^{2}+m^{2}}$
$=\left(\frac{g}{4\pi m^{2}}\right)^{s_{1}}\frac{g}{4\sin\pi s_{1}}\int_{0<{\rm Re}(u)<s_{1}-s}\frac{du}{2\pi i}\frac{\Gamma(u)\Gamma(s-s_{1}+1)\Gamma(-s+s_{1}-u)^{2}}{\Gamma(1-u)\Gamma(s_{1}-s)}\left(\frac{p^{2}}{m^{2}}\right)^{-u}\ .$ (160)
Combining all the above, one obtains the Mellin-Barnes representation for $S(p^{2})(p^{2}+m^{2})^{2}$. The only tricky point is that the $dt$ integral has to be interpreted in the sense of analytical continuation. When $s_{1}$ increases, one must partial integrate more and more times, in order to reach the convergence region. This can lead to additional poles when shifting the $s_{1}$ contour to the right and must be properly taken into account. It is interesting to notice that the $2s-s_{1}=-k$ poles has been canceled by the denominator.
[…]
${\cal H}_{0,k}(p^{2})(p^{2}+m^{2})^{2}$
[…]
$\left(\psi(-k-s_{1}-t+1)+\psi(k+s_{1}+t)-\ln\left(\frac{p^{2}}{m^{2}}\right)+2\gamma_{E}\right)\times\left(\frac{p^{2}}{m^{2}}\right)^{-k-s_{1}-t}\ .$ (161)
Notice that the first letter 00 in ${\cal H}_{0,k}$ denotes the $u=s_{1}-s+0$ pole. Here we would like to shift the $s_{1}$ contour to the right. There are poles at $s_{1}=n\geq 1$ due to the analytical continuation of $s_{1}$ integral. There are another series of poles at $s_{1}=-t+n$ with $n\geq 1$. The non-trivial fact that guarantees the correctness of the calculation is the following: the $s_{1}=-t+n$ poles cancel between ${\cal H}_{0,k}$ and $S_{0,k+n}$, where $S_{0,k+n}$ is the operator contribution of $O_{k+n}$ obtained by picking up the $s=s_{1}-(k+n)$ contribution at the previous step. On the other hand, for the $s_{1}=n$ poles in the ${\cal H}_{0,k}$, if the $dt$ integral is convergent, then one could pickup that pole by partial-integrating $n+1$ times and obtains
${\cal H}_{0,k}\bigg{|}_{s_{1}=n}(p^{2}+m^{2})^{2}$
[…]
×$\times\left(\psi(-k-n-t+1)+\psi(k+n+t)-\ln\left(\frac{p^{2}}{m^{2}}\right)+2\gamma_{E}\right)\left(\frac{p^{2}}{m^{2}}\right)^{-t}\ .$ (162)
Here there are poles on the positive $t$ axis. Although for the $t\geq 1$ poles one could get rid of them by tilting the $t$ contour, the $t=0$ pole can not be handed by this method. Fortunately, one can show that the operator contribution $S_{0,n+k}$ when shifted to $s_{1}=n$ exactly cancels the $t=0$ pole from ${\cal H}_{0,k}|_{s_{1}=n}$. As a result, one obtained the hard contribution as
${\cal H}_{0,k}\bigg{|}_{s_{1}=n}(p^{2}+m^{2})^{2}=\frac{g}{4(p^{2})^{n}}\left(\frac{g}{4\pi}\right)^{n}\left(\frac{m^{2}}{p^{2}}\right)^{k}\frac{d^{n}}{dt^{n}}\bigg{[}\frac{(-1)^{k+n}\,_{2}\tilde{F}_{1}(-k,n-1;k+n+2t;-1)}{k!\sin(2\pi t)\Gamma(-n-2k-2t+1)}$
×$\times\left(\psi(-k-n-t+1)+\psi(k+n+t)-\ln\left(\frac{p^{2}}{m^{2}}\right)+2\gamma_{E}\right)\left(\frac{p^{2}}{m^{2}}\right)^{-t}\bigg{]}_{t=0}.$ (163)
Here the $\frac{d^{n}}{dt^{n}}F(t)|_{t=0}$ is defined as the $n!$ times the $t^{n}$ coefficient of the Taylor expansion of $F(t)$ at $t=0$, even there are negative terms at $t=0$. For example, for $k=0$ (no mass-insertion on the bubbles), the above gives
${\cal H}_{0,0}\bigg{|}_{s_{1}=n}(p^{2}+m^{2})^{2}(p^{2})^{n}$
$=\left(\frac{g}{4\pi}\right)^{n+1}\frac{d^{n}}{dt^{n}}\bigg{[}\left(\frac{m^{2}}{p^{2}}\right)^{t}\left(\psi(-n-t+1)+\psi(n+t)-\ln\left(\frac{p^{2}}{m^{2}}\right)+2\gamma_{E}\right)\bigg{]}_{t=0}\ .$ (164)
This is nothing but the $C^{B}_{n,0}$ term in Eq. (127) that contains factorial growth calculated from IR renormalized bubble-chains, at the natural renormalization scale $\mu^{2}=m^{2}$ that kills “tree-level” operators. As a result, from the Mellin-Barnes one confirms the correctness of the fixed-pointed calculation.
Using Mellin Barnes, one can also verify the cancellation pattern between operators and coefficient functions. In fact, this simply corresponds to the cancellation of $t\geq 1$ poles after picking up $s_{1}\in Z_{\geq 0}$ and implies that the $(1\pm i0)t$ prescriptions introduced separately to the functions $\{{\cal H}_{0,k}\}_{k\geq 0}$ and $\{S_{0,k}\}_{k\geq 1}$ cancel with each-other. For example, the $t=1$ pole of ${\cal H}_{0,0}|_{s_{1}=n}$ cancels with the $t=1$ pole from $S_{0,n+1}\bigg{|}_{s_{1}=n}$. Namely, it is the $g^{n+1}$ contribution to the operator $O_{n+1}$ that cancels the leading factorial growth of $C_{n,0}$, consistent with the cancellation pattern suggested in the previous subsection. More generally, the $t=p$ pole of ${\cal H}_{0,k}\bigg{|}_{s_{1}=n}$ cancels the $t=p$ pole of $S_{0,k+n+p}\bigg{|}_{s_{1}=n}$, thus not only leading renormalon, sub-leading renormalons also cancel between coefficient functions and operators. Moreover, for each $l\geq 1$, the same cancellation pattern applies also to ${\cal H}_{l,k}$ and $S_{l,k}$.
3.4 Leading large $n$ factorial asymptotics
Here we should mention that an important result that is most efficiently obtained from the Mellin-Barnes trick is to derive a closed formula for the leading large-$n$ asymptotics of the full coefficient functions. For this purpose, first notice that the method discussed in the previous subsection allows to obtain the following leading IR-renormalon asymptotics for the contributions with $l$ mass insertions on the $p-k$ line, while with $k$ mass insertions within the bubble-chain
$\left(\frac{m^{2}}{g}\right)^{k+l}C_{n;l,k}(\mu^{2}=m^{2})\bigg{|}_{n\rightarrow\infty,+1}=$
$-\frac{m^{2}}{p^{2}}\frac{(-1)^{k+l}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k+l}\frac{(n!)^{2}(n-k-l)!\Gamma(k-l+n+2)}{k!(l!)^{2}((n-l)!)^{2}}$
×$\times\,_{2}\tilde{F}_{1}(-k,-k-l+n-1;-l+n+2;-1)\ .$ (165)
Corrections to the above are exponentially small in $n$. Using the fact that the largest contribution is clearly supported in the region $n\gg k$, one can perform the large $n$ expansion to obtain
$\left(\frac{m^{2}}{g}\right)^{k}\frac{C_{n;l,k}}{C_{n;l,0}}\bigg{|}_{n\rightarrow\infty,+1}=\frac{(-2)^{k}}{\Gamma(k+1)}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+\frac{(-1)^{k}2^{k-1}}{\Gamma(k-1)}\frac{1}{n}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+{\cal O}\left(\frac{l}{n^{2}}\right)\ .$ (166)
Since the peak of $l$ is around $l=\sqrt{n}$, the terms neglected are subleading in $\frac{1}{n}$. Thus using the fact
$\left(\frac{m^{2}}{g}\right)^{l}C_{n;l,0}(\mu^{2}=m^{2})\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{l}\frac{m^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ ,$ (167)
one obtains after summing over $l$ and $k$, the following leading large $n$ asymptotics originated from the $s=1$ IR renormalons of the $C_{n}$
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2}=m^{2},p^{2})\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\frac{m^{2}}{p^{2}}e^{-\frac{8\pi m^{2}}{g}}n!L_{n}\left(\frac{4\pi m^{2}}{g}\right)\bigg{(}1+\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{2}{n}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (168)
Here $L_{n}$ is the Laguerre polynomial with $L_{n}(0)=1$. The above asymptotics has been checked numerically given Eq. (3.4). For generic $\mu^{2}$, the natural generalization is
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2},p^{2})\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\frac{\mu^{2}}{p^{2}}e^{-\frac{8\pi m_{r}^{2}}{g}}n!L_{n}\left(\frac{4\pi m_{r}^{2}}{g}\right)\bigg{(}1+\left(\frac{4\pi m_{r}^{2}}{g}\right)^{2}\frac{2}{n}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (169)
Clearly, the factorial growth will never be canceled and the power expansion for the coefficient function as well as for the momentum space two-point function will not be convergent. Notice that the over-all factor $e^{-\frac{8\pi m_{r}^{2}}{g}}$ is non-perturbative in the coupling constant $g$.
Similarly, at $\mu^{2}=m^{2}$, from the Mellin Barnes one can derive the following leading non-alternating factorial asymptotics up to subleading renormalons for the operator condensates
$\bigg{\langle}O_{n;k}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty,+1}$ $\rightarrow\frac{(-1)^{k+1}g^{n+1}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k}\frac{(n-k)!\Gamma(k+n)}{\Gamma(k)}$
×$\times\,_{2}\tilde{F}_{1}(1-k,-k+n-1;n+1;-1)\ ,$ (170)
where $O_{n;k}$ denotes the contribution to the operator $O_{n}$ in the $n-k$ bubble diagram with $k\geq 1$. For $k\leq 0$, there is no non-alternating factorial growth, consistent with the UV renormalon interpretation for operators. Combining Eq. (3.4) and Eq. (3.4), one can see that the leading factorial growth of the operators $O_{n+1}$ in the $n-k$ bubble diagram exactly cancel that of $C_{n;0,k}$ with $k$ mass insertions within the $n-k$ bubbles. Now, if there are $l$ mass insertions on the $p-k$ line, one should expect that the $O_{n+1-l;k+1}$ combined with the $(m^{2})^{l}$ term in the coefficient function cancel with $C_{n;l,k}$. Indeed, using results in Appendix. B, one obtains
$\frac{(-1)^{l}m^{2}}{(p^{2})^{n+1}}\frac{g^{n+1}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k+l}\frac{(n-l+k)!^{2}}{l!^{2}(n-l)!^{2}}\times\frac{(n-l-k)!\Gamma(n-l+k+2)}{k!}$
×$\times\,_{2}\tilde{F}_{1}(-k,n-k-l-1;n-l+2;-1)\ ,$ (171)
which exactly cancels Eq. (3.4) multiplying the overall factor $\frac{g^{n+1}}{(p^{2})^{n}}$ for $C_{n}$. This is completely consistent with the general cancellation pattern. Moreover, by summing over $k\geq 0$, the leading non-alternating factorial growth for the operator condensate reads
$\bigg{\langle}O_{n}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty,+1}\rightarrow\frac{m^{2}g^{n}}{(4\pi)^{n}}(n-1)!e^{-\frac{8\pi m^{2}}{g}}\bigg{(}1+\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{2}{n-1}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (172)
Again, the non-alternating factorial asymptotics is exponentially small in $g$. The Eq. (3.4), Eq. (3.4) and Eq. (172) represent one of the major results of this work.
Leading asymptotics from the alternating renormalons can also be calculated in similar manners. For example, at $\mu=m$ one has
$\bigg{\langle}O_{n;-1+k}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty;-1}\rightarrow\frac{(-1)^{n}g^{n+2}}{(4\pi)^{n+2}}\left(\frac{4\pi m^{2}}{g}\right)^{k}\frac{(-k+n+1)!\Gamma(k+n-1)}{k!}$
×$\times\,_{2}\tilde{F}_{1}(-k,n-k;n-1;-1)\ .$ (173)
In particular, for $k=0$ the above exactly reduces to the Eq. (151) generated in the absolutely convergent $n+1$ bubble diagram. Notice that for $k\leq-1$ there are no leading alternating renormalon, while the $s_{1}=-q$ subleading renormalon persists up to $k=-q+1$. Summing over $k$ one obtains
$\bigg{\langle}O_{n}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty;-1}$
$\rightarrow\frac{(-1)^{n}g^{n+2}}{(4\pi)^{n+2}m^{2}}(n+1)!e^{\frac{8\pi m^{2}}{g}}\bigg{(}1-\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{(\frac{g}{\pi m^{2}}-2)}{n}+{\cal O}\left(\frac{1}{g^{4}n^{2}}\right)\bigg{)}\ .$ (174)
Strikingly, the leading alternating factorial growth is exponentially large in $g$. This is not in contradiction with the fact that in the $g\rightarrow 0$ limit there are no factorial growths, since to reach the asymptotic region one needs $n\gg\frac{1}{g^{2}}$, which moves to infinity as $g\rightarrow 0$.
Given the asymptotcis of the operators and the coefficient functions, let’s consider their effects on the power expansion. Then, when projected to any given power $n$ in $(p^{2}+m^{2})^{2}S_{\rm NL}$, in the large $n$ limit one has
$(p^{2}+m^{2})^{2}S_{\rm NL}^{(n)}\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\bigg{(}\left(\frac{\hat{g}}{p^{2}}\right)^{n}(n-1)!L_{n-1}\left(\frac{m^{2}}{\hat{g}}\right)-\left(\frac{\hat{g}}{p^{2}}\right)^{n+1}n!L_{n}\left(\frac{m^{2}}{\hat{g}}\right)\bigg{)}\ ,$ (175)
where $\hat{g}=\frac{g}{4\pi}$. Clearly, the first contribution is due to the operators $O_{n-l},l\geq 0$, while the second contribution is due to the coefficient function $C_{{\rm I},{\rm NL}}^{(n)}$. Given the factorial behavior, the power expansion clearly failed to converge. If one introduces the Borel transforms
$B_{s}[t]=m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\sum_{n=0}^{\infty}t^{n}L_{n}\left(\frac{m^{2}}{\hat{g}}\right)=m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\frac{\exp\left(-\frac{m^{2}}{\hat{g}}\frac{t}{1-t}\right)}{1-t}\ ,$ (176)
$B_{h}[t]=-\frac{\hat{g}}{p^{2}}m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\sum_{n=0}^{\infty}t^{n}(n+1)L_{n+1}\left(\frac{m^{2}}{\hat{g}}\right)=-\frac{\hat{g}}{p^{2}}\frac{d}{dt}B_{s}[t]\ ,$ (177)
then there is a singularity at $t=1$ which is of infinite order. The above implies that the power-expansion for the non-trivial coefficient functions, as well as the sum over non-trivial operators are all Borel non-summable. Despite this divergence, since $B_{s}[t]$ do not grow at infinity, one clearly has for $\frac{\hat{g}}{p^{2}}>0$
[…]
Namely, the full power expansion, although divergent and Borel non-summable in the strict sense, contain no “Borel ambiguity” if only $(1\pm i0)R$ paths are allowed. In fact, this is a general property for sums of the following form $a_{k+q}x^{k+q}-a_{k}x^{k}$, as far as the Borel transform of $\{a_{k}\}$ do not grow too fast at infinity.
4 Comments
Before ending the work, let’s make the following comments.
  1. 1.
    First, we emphasize that due to the off-diagonal cancellation pattern across different powers, the naive power-expansion in $\frac{1}{(p^{2})^{n+2}}$ for the two point function, although free from any ambiguity at any given power, failed to converge. The OPE expansion is also not convergent, unless one redefines all the operators by subtracting out all the factorial growths and adding back to coefficient functions. Even after this procedure, the coefficient function still can not be summed as a power expansion.
  2. 2.
    Second, we should also notice that the renormalization factors in the $\overline{\rm MS}$ scheme for operators do not introduce factorial enhancements, see Eq. (118) for example. This is consistent with the general wisdom that the $\overline{\rm MS}$ scheme renormalization data such as anomalous dimensions do not introduce renormalons Palanques-Mestre and Pascual (1984). The factorial growths of $O_{n}$ in our opinion therefore reflect their genuine properties.
  3. 3.
    We should emphasize that behind the factorial enhancements of operators is a correlation between operator’s derivative numbers and perturbative orders. Indeed, for the operator $O_{n}$ with $n-1$ $\partial^{2}$ inside, it is the bubble-chain diagrams with bubble numbers around $n$ that cause the factorial growth of $O_{n}$ in $n$: for these diagrams, there is a very large power of the logarithms, while a relatively small algebraic power. For bubble numbers $n+k\gg n$, the UV region is suppressed by $\frac{1}{k^{n}}$ and will no-longer cause factorial growth when $k\sim{\cal O}(n)$. This is consistent with the fact that $O_{n}$ themselves have no Borel ambiguities when regarded as perturbative expansions in $\frac{g}{m^{2}}$.
    The situation is different from marginal theory: as least based on lessons from large-$N$ expansion David (1982, 1984, 1986); Beneke et al. (1998); Marino and Miravitllas (2024); Liu (2024), in asymptotically free theories there are ambiguities to define high-dimensional operators, but these ambiguities can be resolved easily in an “one-for-all” manner using a Borel prescription. Once the Borel prescription has been fixed, there are no further factorial enhancements of “high-twist” operators.
  4. 4.
    We should also mention that unlike the renormalons in supper-renormalizable theories, the factorial enhancements observed here are rather subtle and are generated from $\frac{1}{\epsilon}\times\epsilon$ effects. As such, it is less transparent to see its presence using naive $d=2$ running-coupling-type arguments. To see if it works, let’s consider the IR renormalized mass-less bubble chain with
    $F(k^{2})=\frac{1}{4\pi k^{2}}\ln\frac{k^{2}}{\mu^{2}}\ .$ (179)
    Plug it in the diagram with $n$ bubbles, one obtains
    $\frac{-g^{n+1}}{(4\pi)^{n}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{\mu_{0}^{2-d}}{(k^{2})^{n}}\left(\frac{\mu^{2}}{k^{2}}\right)^{s}\frac{n!}{(p-k)^{2}}\bigg{|}_{s^{n}}-\frac{g^{n+1}}{(4\pi)^{n+1}\epsilon}\frac{1}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}\ .$ (180)
    Here $|_{s^{n}}$ denotes the $s^{n}$ power Taylor expansion coefficient at $s=0$. Notice the last term corresponds to the IR subtraction on the $p-k$ line. In this manner, one obtains
    $-\frac{g^{n+1}n!}{(4\pi)^{n+1}(p^{2})^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\exp(\gamma_{E}\epsilon)\Gamma(-\epsilon)\Gamma(-\epsilon-n-s+1)\Gamma(\epsilon+n+s)}{\Gamma(n+s)\Gamma(-2\epsilon-n-s+1)}\bigg{|}_{s^{n}}$
    $-\frac{g^{n+1}}{(4\pi)^{n+1}\epsilon}\frac{1}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}$
    $\rightarrow\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}n!\left(-l_{\mu}+\psi(-n-s+1)+\psi(n+s)+2\gamma_{E}\right)\bigg{|}_{s^{n}}\ .$ (181)
    Again, the $\ln p$ in the $\frac{1}{\epsilon}$ is removed by the IR subtraction on $p-k$, but one is left with a pole at $s=0$. This pole is related to the operator renormalization for $O_{n}$ that contains no dependency on $p$ and can not be calculated accurately in this method. Neglecting this pole, one obtains exactly the Eq. (127) for $C_{n,0}^{B}$ that contains the factorial growth in $n$. Therefore, similar to the 4D bubble chain in QCD Beneke (1999), the naive bubble chain approximation here also produces the correct large $n$ factorial behavior, up to a scheme dependent constant part that is related to operator renormalization and contains no factorial growth in $\overline{\rm MS}$-like schemes.
    The factorial growth in fact allows a simpler explanation: if one consider the integrand
    $\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\ll\mu^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{n}}\ln^{n}\frac{k^{2}}{\mu^{2}}\frac{1}{(p-k)^{2}}\ ,$ (182)
    then for small-$k$, although it has a high degree of IR divergence, the large-power of $k^{2}$ seems to have balanced out the large power of the logarithm and no-factorial behavior seems possible. However, if one Taylor expand the integrand to the first power where the integral converges in the IR, then one has
    $\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\ll\mu^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{n}}\ln^{n}\frac{k^{2}}{\mu^{2}}\frac{(k^{2})^{n}}{(p^{2})^{n+1}}\rightarrow-\frac{g^{n+1}}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}\frac{n!}{(p^{2})^{n}}\ ,$ (183)
    which is just the leading non-alternating factorial growth in Eq. (128) after multiplying the overall factor $\frac{g^{n+1}}{(p^{2})^{n}}$. From the above, it is clear that the factorial growth in the coefficient function is produced in the region where $k^{2}\ll\mu^{2}$ in the renormalized bubbles. Naively, in the IR subtraction, $\mu$ serves as the separation point between UV and IR and coefficient functions should not go the momentum region below than $\mu$. The point is, here the subtraction is local in nature (they are just the UV poles for operators) and can not see the momenta within the massless bubble. Even when the $\mu$ is actually larger than the $k^{2}$ of the bubble, one still subtracts up to the point $\mu$ which is actually much harder than the massless bubble. This is completely consistent with the interpretation of the observed factorial growth as an IR renormalon.
  5. 5.
    Finally, we must emphasize that the factorial enhancement is only $n!$. This is sufficient to make the momentum space OPE divergent. However, in the coordinate space expansion, there is always an additional $\frac{1}{(n!)^{2}}$ suppression factor for the $(z^{2})^{n}$ power, thus a single $n!$ can not affect the convergence of coordinate space OPE or power expansion at large $N$. Moreover, beyond the $\frac{1}{N}$ expansion, there can be factorial growth in the numbers of Feynman-diagrams. But as far as the number of Feynman diagrams at order $g^{n}$ is bounded by $n!\times C^{n}$, the factorial enhancement for the $n$-the power in the coefficient functions would at worst lead to $\frac{n!\times n!}{(n!)^{2}}\tilde{C}^{n}(gz^{2})^{n\pm 1}$, which still has a finite convergence radius. This is consistent with the statements in Lukyanov and Zamolodchikov (1997).
Appendix A Expansion of the sunrise diagram
We consider the $g^{2}$ order sunset diagram for the scalar two point function (this is the only diagram at order $g^{2}$)
$S(z^{2}m^{2})=\frac{g^{2}}{6}\int\frac{d^{2}p}{(2\pi)^{2}}\frac{e^{ip\cdot z}}{(p^{2}+m^{2})^{2}}\Sigma(p^{2})\ ,$ (184)
$\Sigma(p^{2})=\int\frac{d^{2}kd^{2}l}{(2\pi)^{4}}\frac{1}{(k^{2}+m^{2})(l^{2}+m^{2})((p-k-l)^{2}+m^{2})}\ .$ (185)
In the appendix without otherwise mentioning we use the letter $S$ only to denote the order $g^{2}$ contribution to the full two-point function. The simplest way of expanding $S(z^{2}m^{2})$ is to use twice the equation of motion to obtain
$(-\partial^{2}+m^{2})^{2}S(z^{2}m^{2})=\frac{g^{2}}{6}G_{0}^{3}(z^{2}m^{2})\ ,$ (186)
Given the first two terms in the small $z^{2}$ expansion, coefficients for higher order terms can all be determined from this equation. For example, we can write
$S(z^{2}m^{2})=\frac{g^{2}}{6m^{4}}\frac{1}{(4\pi)^{3}}\bigg{(}\alpha+\beta\frac{z^{2}m^{2}}{4}+\left(\frac{z^{2}m^{2}}{4}\right)^{2}\left(aL_{z}^{3}+bL_{z}^{2}+cL_{z}+d\right)\bigg{)}\ .$ (187)
Using $\partial^{2}=\frac{d^{2}}{dz^{2}}+\frac{1}{z}\frac{d}{dz}$, and $G_{0}=-\frac{1}{4\pi}L_{z}+{\cal O}(z^{2}m^{2}L_{z})$, by matching all the logarithms and constants at the order $z^{0}$, one can solve
$4a=-1\ ,4b+36a=4c+78a+24b=4d+36a+26b+12c+\alpha-2\beta=0\ .$ (188)
Thus
$a=-\frac{1}{4},\ b=\frac{9}{4},\ c=-\frac{69}{8},\ d=\frac{27}{2}-\frac{\alpha}{4}+\frac{\beta}{2}\ .$ (189)
By calculating the diagrams for the $\phi^{2}$, $\frac{z^{2}}{4}\phi\partial^{2}\phi$ contributions as in Sec. 2, one can also determine $\alpha=\frac{7\zeta_{3}}{4}$, $\beta=-\frac{21\zeta_{3}}{4}$, thus $d=-\frac{7\zeta_{3}}{16}-\frac{42\zeta_{3}}{16}=-\frac{49\zeta_{3}}{16}$. However, for illustration purpose, here we present the direct approach based on Mellin-Barnes.
The first step is to introduce the Schwinger parameters for the self energy
$\Sigma(p^{2})=\frac{1}{(4\pi)^{2}}\int\frac{d\alpha_{1}d\alpha_{2}d\alpha_{3}}{\alpha_{1}\alpha_{2}+\alpha_{1}\alpha_{3}+\alpha_{2}\alpha_{3}}\exp\bigg{[}-p^{2}\frac{\alpha_{1}\alpha_{2}\alpha_{3}}{\alpha_{1}\alpha_{2}+\alpha_{1}\alpha_{3}+\alpha_{2}\alpha_{3}}-m^{2}\alpha_{123}\bigg{]}\ .$ (190)
Now, we perform Mellin transform with respect to $\frac{p^{2}}{m^{2}}$
$\Sigma(p^{2})=\frac{1}{(4\pi)^{2}m^{2}}\int\frac{ds}{2\pi i}\left(\frac{p^{2}}{m^{2}}\right)^{-s}{\cal M}(s)\ ,$ (191)
${\cal M}(s)=\Gamma(s)\int_{0}^{\infty}\frac{(\alpha_{1}\alpha_{2}\alpha_{3})^{-s}d\alpha_{1}d\alpha_{2}d\alpha_{3}}{(\alpha_{1}\alpha_{2}+\alpha_{1}\alpha_{3}+\alpha_{2}\alpha_{3})^{1-s}}\exp\bigg{[}-(\alpha_{1}+\alpha_{2}+\alpha_{3})\bigg{]}\ .$ (192)
Here $10<{\rm Re}(s)<$ 1. It is a meromorphic function with poles at integers. Right to the convergence strip $10<{\rm Re}(s)<$ 1, it contains an overall $s=1$ pole corresponding to the $\alpha_{1}\sim\alpha_{2}\sim\alpha_{3}\rightarrow 0$ region, and also contains poles from the sub-regions $\alpha_{1}\gg\alpha_{2}\sim\alpha_{3}$ or $\alpha_{1}\sim\alpha_{2}\gg\alpha_{3}$. Using $\alpha_{1}=\beta_{0}$, $\alpha_{2}=\beta_{0}\beta_{1}$, $\alpha_{3}=\beta_{0}\beta_{1}\beta_{2}$, one can de-singularize all these poles and see the analyticity structure directly
[…]
From the above, it is clear that all the singularities have been factorized to the overall factor $(\beta_{1}\beta_{2})^{-s}$, while the rest of the integrand are never singular. Moreover, by partial integrating, ${\cal M}(s)$ becomes a meromorphic function in $s$ with poles at $s=n\in Z$. Furthermore, the ${\cal M}(s)$ decays exponentially as ${\cal O}(e^{-\pi|{\rm Im}(s)|})$ in the vertical strip, while grows in ${\rm Re}(s)$ generically no faster than ${\cal O}\left(9^{{\rm Re}(s)}\right)$. This is consistent with the general pattern of Mellin transform for momentum space correlation function in fixed order perturbation theory.
Although the parametric representation can be useful to establish general properties of ${\cal M}(s)$, for calculation purposes this representation is rather inconvenient. Instead of using the parametric representation, the expansion of ${\cal M}(s)$ around its poles can be systematically calculated using another layer of Mellin Barnes. One can first introduce Mellin representation for $\tilde{\Sigma}(p^{2})=\int\frac{d^{2}k}{(k^{2}+m^{2})((p-k)^{2}+m^{2})}$, then obtain that for $\Sigma(p^{2})=\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\tilde{\Sigma}(k^{2})}{(p-k)^{2}+m^{2}}$. This leads to the Barnes representation
${\cal M}(s)=\int\frac{dz}{2\pi i}\frac{\Gamma(1-z)^{4}\Gamma(s)\Gamma(z-s)^{2}}{\Gamma(2-2z)\Gamma(1-s)}\ ,$ (194)
where $10<{\rm Re}(z)<$ 1 and $10<{\rm Re}(s)<{\rm Re}(z)<$ 1. Thus, by shifting the contour of $z$ to the left and pick up the poles at $z=s-k,k\geq 0$, one obtains the following absolutely-convergent expansion
${\cal M}(s)=\sum_{k=0}^{\infty}{\cal M}_{k}(s)\ ,$ (195)
${\cal M}_{k}(s)=\frac{2\Gamma(s)\left(-2H_{k-s}+H_{2k+1-2s}+H_{k}\right)\Gamma(k+1-s)^{4}}{(k!)^{2}\Gamma(1-s)\Gamma(2k+2-2s)}\ .$ (196)
Notice that ${\cal M}_{k\geq n}(s)$ vanishes at $s=n\geq 1$, thus only ${\cal M}_{0\leq k\leq n-1}$ are required to calculate the expansion of ${\cal M}(s)$ up to constant terms at $s=n,n\geq 1$. In particular, the expansion of ${\cal M}(s)$ for $s=1,2,3$ up to constant term read
${\cal M}(s)\bigg{|}_{s\rightarrow 1}=-\frac{6}{(s-1)^{3}}+4\zeta_{3}\ ,$ (197)
${\cal M}(s)\bigg{|}_{s\rightarrow 2}\rightarrow-\frac{6}{s-2}+\frac{24}{(s-2)^{2}}+\frac{18}{(s-2)^{3}}-12(\zeta_{3}+1)\ ,$ (198)
${\cal M}(s)\bigg{|}_{s\rightarrow 3}\rightarrow\frac{3}{2(s-3)}-\frac{156}{(s-3)^{2}}-\frac{90}{(s-3)^{3}}+\left(60\zeta_{3}+\frac{183}{2}\right)\ .$ (199)
And poles at $s=n\geq 4$ can be calculated systematically.
Given the Mellin representation of $\Sigma(p^{2})$, it is straightforward to obtain the Mellins representation of the full two-point function using
[…]
$=\frac{1}{4\pi m^{2}}\int\frac{du}{2\pi i}\left(\frac{z^{2}m^{2}}{4}\right)^{-u}\frac{\Gamma(u)\Gamma(1-s-u)\Gamma(1+s+u)}{\Gamma(1-u)}\ .$ (200)
Here the contour is $0<{\rm Re}(u)<1-{\rm Re}(s)$. Thus
$S(z^{2}m^{2})=\frac{g^{2}}{6}\frac{1}{(4\pi)^{3}m^{4}}\int\frac{ds}{2\pi i}\int\frac{du}{2\pi}\frac{\Gamma(u)\Gamma(1-s-u)\Gamma(1+s+u){\cal M}(s)}{\Gamma(1-u)}\left(\frac{z^{2}m^{2}}{4}\right)^{-u}\ .$ (201)
Now, one can obtain the expansion of $S(z^{2})$ by shift the $u$ contours to the left, and then shift the $s$ contours to the right. The poles at $u=-l\leq 0$ will be the “soft” contributions at powers $(z^{2})^{l}$, while the pole at $u=-l-s,l\geq 1$ then $s=n\geq 1$ will contain the “hard contributions” at powers $(z^{2})^{n+l}$, namely, contribution from the regions where at least one line is hard.
We first consider the hard contributions. For our purpose one needs
${S}_{u=-1-s}(z^{2}m^{2})=\frac{g^{2}}{6}\frac{1}{(4\pi)^{3}m^{4}}\int\frac{ds}{2\pi i}\frac{{\cal M}(s)\Gamma(-1-s)}{\Gamma(s+2)}\left(\frac{z^{2}m^{2}}{4}\right)^{s+1}\ .$ (202)
Using the expansion of ${\cal M}(s)$ at $s=1$ one obtains
$\Sigma(p^{2})\bigg{|}_{p^{2}\rightarrow\infty}\rightarrow\frac{3}{(4\pi)^{2}p^{2}}\ln^{2}\frac{p^{2}}{m^{2}}+{\cal O}\left(\frac{m^{2}}{p^{4}}\right)\ ,$ (203)
$S_{u=-1-s}\bigg{|}_{z^{2}\rightarrow 0}\rightarrow\frac{1}{6(4\pi)^{3}}\bigg{(}\frac{z^{2}g}{4}\bigg{)}^{2}\bigg{(}-\frac{L_{z}^{3}}{4}+\frac{9L_{z}^{2}}{4}-\frac{69L_{z}}{8}+\frac{27}{2}\bigg{)}\ ,$ (204)
with $L_{z}=\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}$. Moreover, using the expansion of ${\cal M}(s)$ at $s=2,3$ allows the expansion for hard parts to be performed up to $z^{8}$.
We then consider purely soft contributions from $u=0,-1,-2,...$ poles. Up to $z^{4}$, there are three numbers which are due to $\langle\phi^{2}\rangle_{g^{2}}$, $\langle\phi\partial^{2}\phi\rangle_{g^{2}}$ and $\langle\phi\partial^{4}\phi\rangle_{g^{2}}$. These numbers only affect expansion in coordinate space. The expansion for the $u=-n$ poles are
$S_{u\in Z_{\leq 0}}(z^{2}m^{2})=\frac{g^{2}}{6(m^{2})^{2}(4\pi)^{3}}\sum_{n=0}^{\infty}\left(\frac{z^{2}m^{2}}{4}\right)^{n}s_{n}\ ,$ (205)
$s_{n}=\frac{1}{(n!)^{2}}\int\frac{ds}{2\pi}\frac{\pi(s-n)}{\sin\pi s}{\cal M}(s)\ .$ (206)
There are two ways of evaluating the integrals. The first and most rigorous method, is to use again the representation
${\cal M}(s)=\int\frac{dz}{2\pi i}\frac{\Gamma(1-z)^{4}\Gamma(s)\Gamma(z-s)^{2}}{\Gamma(2-2z)\Gamma(1-s)}\ ,$ (207)
where $10<{\rm Re}(z)<$ 1 and $10<{\rm Re}(s)<{\rm Re}(z)<$ 1. Thus, one can perform first the $ds$ integral within $dz$ using the first Barnes lemma, which leads to
$a=\int\frac{ds}{2\pi i}\frac{\pi s}{\sin\pi s}{\cal M}(s)=\int\frac{dz}{2\pi i}\frac{\pi^{3}z\cot(\pi z)\csc^{2}(\pi z)}{1-2z}\ ,$ (208)
$b=\int\frac{ds}{2\pi i}\frac{\pi}{\sin\pi s}{\cal M}(s)=\int\frac{dz}{2\pi i}\frac{2\pi^{3}\cot(\pi z)\csc^{2}(\pi z)}{1-2z}\ .$ (209)
The above can be evaluated by shifting the contour to the right
$a=\sum_{k=1}^{\infty}\frac{2}{(2k-1)^{3}}=\frac{7}{4}\zeta_{3}\ ,$ (210)
$b=\sum_{k=1}^{\infty}\frac{8}{(2k-1)^{3}}=7\zeta_{3}\ .$ (211)
This determines the soft part of the expansion as
$S_{u\in Z_{\leq 0}}(z^{2}m^{2})=\frac{g^{2}}{6m^{4}}\frac{7\zeta_{3}}{(4\pi)^{3}}\sum_{n=0}^{\infty}\frac{1}{(n!)^{2}}\left(\frac{1}{4}-n\right)\left(\frac{z^{2}m^{2}}{4}\right)^{n}\ .$ (212)
In particular, the contribution at $(z^{2}m^{2})^{4}$ reads
$S_{u=-2}=-\frac{g^{2}z^{4}}{6}\frac{49\zeta_{3}}{256(4\pi)^{3}}\ .$ (213)
Moreover, the soft part converges absolutely for any $z^{2}$ and can be summed into Bessel $I$ function. Notice that the soft contribution grows exponentially at large $z$. Both soft and hard contributions are required to guarantee exponential decay.
To summarize, the full expansion of the two point function up to power $z^{4}$ and ${\cal O}(g^{2})$ reads
$S(z^{2}m^{2},\frac{g}{m^{2}})=$
$-\frac{1}{4\pi}L_{z}+\frac{7\zeta_{3}}{24(4\pi)^{3}}\frac{g^{2}}{m^{4}}$
$+\bigg{(}\frac{2-L_{z}}{4\pi}-\frac{7\zeta_{3}}{8(4\pi)^{3}}\frac{g^{2}}{m^{4}}\bigg{)}\frac{z^{2}m^{2}}{4}$
$+\bigg{(}\frac{3-L_{z}}{16\pi}+\frac{g^{2}}{6(4\pi)^{3}m^{4}}\left(-\frac{L_{z}^{3}}{4}+\frac{9L_{z}^{2}}{4}-\frac{69L_{z}}{8}+\frac{27}{2}-\frac{49\zeta_{3}}{16}\right)\bigg{)}\left(\frac{z^{2}m^{2}}{4}\right)^{2}\ .$ (214)
Here we have added the expansion of the tree-level diagram.
Appendix B Angular averages
In this appendix we derive the angular averages for the small-$k$ expansion of the tree-level propagator $\frac{1}{(p-k)^{2}+m^{2}}$ in an arbitrary $d$-dimension.
The simplest way to perform this expansion is to use the following trick. Consider
$I(z,k)=\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{(p-k)^{2}+m^{2}}=\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{p^{2}+m^{2}}e^{ik\cdot z}\ .$ (215)
Clearly, in order to obtain the small $k$ expansion from the region where $p$ is hard, it is sufficient to keep only the hard contribution for the coordinate-space two-point function and then Fourier-transform back. For this purpose, one can write
$\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{p^{2}+m^{2}}\bigg{|}_{\rm hard}=\sum_{l=0}^{\infty}\int\frac{d^{D}p}{(2\pi)^{D}}\frac{(-1)^{l}m^{2l}e^{ip\cdot z}}{(p^{2})^{l+1}}$
$=\sum_{l=0}^{\infty}(-1)^{l}m^{2l}\frac{\Gamma\left(\frac{D}{2}-l-1\right)}{(4\pi)^{\frac{D}{2}}\Gamma(l+1)}\left(\frac{z^{2}}{4}\right)^{1+l-\frac{d}{2}}\ .$ (216)
Now, for the angular average of $e^{ik\cdot z}$, one has
[…]
Given the above, by Fourier transforming back from the coordinate space to the momentum space again, one obtains after angular average over $k$
$\frac{1}{(p-k)^{2}+m^{2}}=\sum_{n=0}^{\infty}\frac{k^{2n}}{(p^{2})^{n+1}}{\cal C}_{n}\left(\frac{m^{2}}{p^{2}}\right)\ ,$ (218)
${\cal C}_{n}\left(\frac{m^{2}}{p^{2}}\right)=\sum_{l=0}^{\infty}\left(\frac{m^{2}}{p^{2}}\right)^{l}\frac{\Gamma\left(1-\epsilon\right)(-1)^{l+n}\Gamma\left(-l-\epsilon\right)(l+n)!}{l!n!\Gamma\left(1-\epsilon+n\right)\Gamma\left(-l-n-\epsilon\right)}$
$=\sum_{l=0}^{\infty}\left(\frac{m^{2}}{p^{2}}\right)^{l}(-1)^{l}\frac{(l+n)!}{l!n!}\frac{\Gamma(1-\epsilon)}{\Gamma(l+1+\epsilon)}\frac{\Gamma(n+l+1+\epsilon)}{\Gamma(n+1-\epsilon)}$
$=\frac{\Gamma(1-\epsilon)\Gamma(n+1+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n+1-\epsilon)}\,_{2}F_{1}\left(n+1,n+1+\epsilon;1+\epsilon;-\frac{m^{2}}{p^{2}}\right)\ .$ (219)
In the large $n$ limit, $C_{n}$ is absent of factorial growth. For example, at $p^{2}=m^{2}$ and $\epsilon=0$, one has
${\cal C}_{n}\bigg{|}_{p^{2}=m^{2},D=2}=\frac{\Gamma\left(\frac{n}{2}+\frac{3}{2}\right)}{\Gamma(n+2)\Gamma\left(-\frac{1}{2}-\frac{n}{2}\right)}\ .$ (220)
It vanishes at $n=2k+1$, while for $n=2k$ it decays exponentially as ${\cal O}\left(2^{-n}\right)$.